### v1

# Stochastic Online Metric Matching: Adversarial is no Harder than Stochastic

Abstract
We study the stochastic online metric matching problem. In this problem, $m$ servers and $n$ requests are located in a metric space, where all servers are available upfront and requests arrive one at a time. In particular, servers are adversarially chosen, and requests are independently drawn from a known distribution. Upon the arrival of a new request, it needs to be immediately and irrevocably matched to a free server, resulting in a cost of their distance. The objective is to minimize the total matching cost.
In this paper, we show that the problem can be reduced to a more accessible setting where both servers and requests are drawn from the same distribution by incurring a moderate cost. Combining our reduction with previous techniques, for $[0,1]^{d}$ with various choices of distributions, we achieve improved competitive ratios and nearly optimal regrets in both balanced and unbalanced markets. In particular, we give $O(1)$-competitive algorithms for $d\geq 3$ in both balanced and unbalanced markets with smooth distributions. Our algorithms improve on the $O((\log\log\log n)^{2})$ competitive ratio of [GGPW19] for balanced markets in various regimes, and provide the first positive results for unbalanced markets.
1 Introduction
Online matching forms a central topic of online algorithms and has received extensive attention due to the surge of online matching markets in recent years. Some of its prominent applications include online advertising, job search, and dating, where the constant growth of online advertising platforms has particularly spurred numerous research from both algorithmic and economic perspectives. A seminal paper [KVV90] proposes the online bipartite matching problem and presents optimal deterministic and randomized algorithms. Since then, elegant algorithmic results have been devised to accommodate more diverse environments. See the recent survey by [Meh13] and the references therein for a more comprehensive introduction.
In recent years, another online matching paradigm, online (bipartite) minimum cost matching, has emerged. Motivated by the application of ride-hailing, in this problem, $m$ nodes from one side of the market are called servers and are fully known upfront. On the other side of the market, which is often referred to as the request side, $n$ nodes arrive sequentially. Upon the arrival of a request $i$, it has to be immediately and irrevocably assigned to an available server $j$, incurring a cost $\delta(i,j)$, and both request $i$ and server $j$ are removed [^1] We always assume $m\geq n$ to guarantee that every request can be served.. The objective is to minimize the total matching cost.
Due to the importance of application to ride-hailing, it is reasonable to assume that both servers and requests are located in the same metric space, and the cost incurred by matching a server and a request is measured by the distance between them. A majority of literature on online metric matching focuses on the adversarial model [PS23, Rag18, MNP06], where the locations of all servers and requests are adversarially chosen, or the weaker random order model [Rag16], which assumes that the order of adversarially chosen requests is randomly permuted. See Section 1.2 for a more comprehensive literature review in these settings.
The online matching model with adversarial online arrivals is usually deemed too pessimistic since practical instances generally exhibit a more stochastic nature. To model a practical and hence more tractable scenario, the stochastic online metric matching problem, where the requests are independently drawn from a known probability distribution, has attracted considerable attention in the past few years [GGPW19, CKKZ23, BFP23]. The most basic model assumes balanced markets, which contain an equal number of servers and requests. For the Euclidean space $[0,1]^{d}$, where $d$ is a constant not scaling with $n$, [Kan22] proposes the Hierarchical Greedy algorithm with a constant competitive ratio when all servers and requests are uniformly distributed, and [CKKZ23] achieves nearly optimal regret guarantees when servers and requests follow different distributions. [BFP23] analyzes the performance of the Greedy algorithm, whose prominence is empirically validated, and shows that Greedy is $O(1)$-competitive when servers and requests are uniformly distributed in $[0,1]$.
[AALS22] notices that the potential value of excess servers has been overlooked by the literature and considers unbalanced markets, which possess at most a linear amount of excess servers. When servers and requests are uniformly distributed in $[0,1]$, they show that the total matching cost of Greedy can be substantially improved by transitioning from balanced markets to unbalanced markets, and their results imply that Greedy is $O(\log^{3}n)$-competitive. [BFP23] subsequently pinpoints the competitive ratio of Greedy to be $O(1)$ under the same setting. Moreover, Hierarchical Greedy also admits a constant competitive ratio in unbalanced markets when servers and requests are uniformly distributed in $[0,1]^{d}$ for all $d\geq 1$ [Kan22].
In contrast to the fruitful achievements when both servers and requests are stochastic, it appears to be more challenging to establish positive results for adversarial servers and stochastic requests. Specifically, in balanced markets, [GGPW19] gives an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, which further achieves a constant competitive ratio for tree metrics. Moreover, in balanced markets with requests uniformly distributed in $[0,1]$, [BFP23] shows that Greedy is $\Theta(\log n)$-competitive and Hierarchical Greedy is $\Omega(n^{1/4})$-competitive.
1.1 Our Results
As our main contribution, assuming requests are stochastic, we present a reduction from the setting with adversarial servers to that with stochastic servers: If an algorithm ${\mathcal{A}}$ exists for balanced markets with both servers and requests randomly drawn from the same distribution $\mathbb{D}$, we can use ${\mathcal{A}}$ in a black-box manner to derive another algorithm ${\mathcal{A}}^{\prime}$ for unbalanced markets with servers adversarially chosen and requests randomly drawn from $\mathbb{D}$ (Theorem 1). Moreover, if the markets are balanced, which can be viewed as a special case of unbalanced markets, ${\mathcal{A}}^{\prime}$ admits the same competitive ratio as ${\mathcal{A}}$ up to a constant factor. Hence, by utilizing the recent progress on the setting with stochastic servers and requests [Kan22, CKKZ23], and applying our reduction techniques, we then describe the following results.
Balanced Markets Unbalanced Markets
Uniform Dis. Smooth Dis. General Dis. Smooth Dis. General Dis.
$d=1$ $O(1)$ $O(1)$ $O(1)$ $O(\log n)$ $O(\log n)$
$d=2$ $O(1)$ ($\omega(1)$) $\omega(1)$ $\omega(1)$ $O(\sqrt{\log n})$ $O(\log n)$
$d\geq 3$ $O(1)$ ($\omega(1)$) $O(1)$ ($\omega(1)$) $\omega(1)$ $O(1)$ $O(\log n)$
Table 1: Competitive ratios for the Euclidean space $[0,1]^{d}$ with adversarial servers and stochastic requests. All previous state-of-the-art results are given in black fonts, where we use $\omega(1)$ to denote $O((\log\log\log n)^{2})$, and all new results achieved in this paper are marked in red. In particular, for balanced markets with general distributions, the competitive ratio of $O(1)$ for $d=1$ is obtained by [GGPW19] and Theorem 9, which implies the two bounds on its left, and the competitive ratio of $O((\log\log\log n)^{2})$ for $d\geq 2$ is achieved by [GGPW19], which implies all $O((\log\log\log n)^{2})$ bounds on its left.
Multiplicative bound (competitive ratio).
We first introduce the competitive ratio results for the Euclidean space $[0,1]^{d}$ achieved by combining our reduction with algorithms of prior work, which are summarized in Table 1. Recall that the competitive ratio of an algorithm is defined as the worst ratio between the cost incurred by the algorithm and the optimal offline cost, and can be viewed as a multiplicative guarantee. We start by showing that combining our reduction with the Hierarchical Greedy algorithm of [Kan22] leads to an $O(1)$-competitive algorithm in balanced markets for all $d\geq 1$ with adversarial servers and uniformly distributed requests (Theorem 2).
Next, we describe our results in balanced markets with servers adversarially chosen and requests randomly drawn from a smooth distribution, where a distribution is smooth if it admits a density function upper bounded by a constant. For the algorithm ${\mathcal{A}}$ required by the reduction, we adopt the $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ ($\mathsf{SOAR}$) algorithm of [CKKZ23], which is originally used to give nearly optimal regrets in balanced markets when servers and requests follow different distributions. We show that this leads to an $O(1)$-competitive algorithm for $d\geq 3$ (Theorem 3).
Finally, we present our results in unbalanced markets with adversarial servers and stochastic requests. When requests are randomly drawn from a smooth distribution, by applying the same approach as for balanced markets, the resulting algorithm admits a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$ (Theorem 4). To improve the undesirable competitive ratio for $d=1$, we show that the $O(\log n)$ competitive ratio of the algorithm proposed by [GL12], which holds in balanced markets when both servers and requests are adversarially located in a doubling metric [^2] A metric space is said to be doubling if a doubling constant $M>0$ exists such that for all $r>0$, every ball of radius $r$ can be covered by at most $M$ balls of radius $r/2$. In particular, the Euclidean space $[0,1]^{d}$ is a doubling metric for constant $d$., is preserved in unbalanced markets (Corollary 2).
Additive bound (regret).
We turn to our regret results for the Euclidean space $[0,1]^{d}$. Since the regret of an algorithm is defined as the worst difference between the average cost incurred by the algorithm and the average optimal offline cost, it can be viewed as an additive guarantee. Recall that in balanced markets with servers and requests following different but arbitrary distributions, [CKKZ23] shows that the $\mathsf{SOAR}$ algorithm admits a regret of $O(n^{-\frac{1}{2}})$ for $d=1$, $O(n^{-\frac{1}{2}}\log n)$ for $d=2$, and $O(n^{-\frac{1}{d}})$ for $d\geq 3$, which is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$. By combining our reduction with the $\mathsf{SOAR}$ algorithm, we show that the same regrets are achievable even in unbalanced markets with servers adversarially chosen and requests randomly drawn from an arbitrary distribution (Theorem 5). Since balanced markets form a special case of unbalanced markets and our setting with adversarial servers is more general, the lower bounds by [CKKZ23] continue to hold in our setting, and hence, our regret guarantees are nearly optimal.
Additional results.
As alluded to previously, we generalize the $O(\log n)$ competitive ratio of the algorithm by [GL12] for doubling metrics with adversarial servers and requests from balanced markets to unbalanced markets. This is achieved based on the observation that every greedy algorithm with a tie-breaking rule satisfying certain properties admits the same competitive ratio in balanced and unbalanced markets (Theorem 6). As a by-product, we similarly show that the $O(\log^{3}n)$ competitive ratio of the algorithm by [MNP06] in balanced markets with servers and requests adversarially located in an arbitrary metric is retained in unbalanced markets (Corollary 1). We remark here that all prior results for adversarial servers and requests are established in balanced markets, and it is not immediate whether they can be carried over to unbalanced markets.
Moreover, in Appendix B, we extend our reduction to the case where cost functions only approximately satisfy the triangle inequality. Specifically, we say that a cost function $\delta$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z$, we have $\delta(x,z)\leq\eta\cdot(\delta(x,y)+\delta(y,z))$. To showcase the generality of our approach, we show that combining our generalized reduction with the $\mathsf{SOAR}$ algorithm leads to positive results for $[0,1]^{d}$ with cost function $\left\|{\cdot}\right\|_{2}^{p}$ for some $p\geq 1$.
Finally, in Appendix C, we show that combining our reduction with the $\mathsf{SOAR}$ algorithm gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions (Theorem 9), recovering the result by [GGPW19].
1.2 Related Work
For general metrics with adversarial servers and requests, [KP93], [KMV94], and [Rag16] respectively give $(2n-1)$-competitive deterministic algorithms, which is also shown to be tight for deterministic algorithms. For randomized algorithms, an $O(\log^{3}n)$-competitive algorithm is proposed by [MNP06]; it is subsequently improved by [BBGN14] to $O(\log^{2}n)$, which remains the best upper bound. In the random-order model, the deterministic algorithm of [Rag16] admits a competitive ratio of $O(\log n)$, which is optimal even for randomized algorithms, and [GK19] shows that the competitive ratio of Greedy lies between $\Omega(n^{0.22})$ and $O(n)$.
For the line metric with adversarial servers and requests, [KP93] shows that the competitive ratio of Greedy is at least $12^{n}-$ 1, and a deterministic algorithm with a sublinear competitive ratio is known [ABN+19]. Furthermore, [NR17] presents an $O(\log^{2}n)$-competitive deterministic algorithm, which is later shown to be $O(\log n)$-competitive [Rag18]. Moving to randomized algorithms, [GL12] proposes an $O(\log n)$-competitive algorithm for the more general doubling metrics. For lower bounds, [FHK05] shows that no deterministic algorithm can achieve a competitive ratio strictly better than $9.001$; it is then improved to $\Omega(\sqrt{\log n})$, which holds even for randomized algorithms [PS23].
Various papers study other aspects of stochastic online metric matching. [TTC94] shows that Greedy admits a competitive ratio of $O(\sqrt{n})$ on the unit disk. In addition, [SNL+23] characterizes the cost of Greedy in unbalanced markets with servers and requests uniformly distributed in $[0,1]$ using a partial differential equation approach.
2 Preliminaries
In our model, $n$ requests and $m$ servers are located in a metric space $(X,\delta)$, where $n\leq m$ and $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ is a cost function satisfying the triangle inequality. Denote the locations of servers as $S=\{s_{1},\ldots,s_{m}\}$ and the locations of requests as $R=\{r_{1},\ldots,r_{n}\}$. The locations of servers are adversarially chosen and all available at the beginning. The requests arrive one by one, and their locations are independently drawn from a known distribution $\mathbb{D}$ over $X$. Upon the arrival of a new request, we need to match it to an available server immediately and irrevocably, and then both the request and the matched server leave the system. The cost of matching a server $s$ and a request $r$ is $\delta(s,r)$, and we aim to minimize the total matching cost.
For all algorithm ${\mathcal{A}}$ and set of servers $S$ with $|S|\geq n$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,n)$ as the expected cost of ${\mathcal{A}}$ when the initial set of servers is $S$ and $n$ requests are randomly drawn from $\mathbb{D}$, and denote $\mathrm{cost}_{{\mathcal{A}}}(n)$ as the expected cost of ${\mathcal{A}}$ when $|S|=|R|=n$ and all servers and requests are randomly drawn from $\mathbb{D}$. Given a matching $M$ between two sets $S$ and $R$, for each matched element $s\in S$ (resp. $r\in R$), we use $M(s)$ (resp. $M(r)$) to denote the element in $R$ (resp. $S$) that $s$ (resp. $r$) is matched to.
We define the smoothness property of distributions as follows, which requires distributions to admit an upper bounded density function.
Definition 1 ($\beta$-smoothness).
We say that a distribution $\mathbb{D}$ over $X$ is $\beta$-smooth if $\mathbb{D}$ admits a density function $\mu_{\mathbb{D}}$ such that $\mu_{\mathbb{D}}(x)\leq\beta$ for all $x\in X$.
2.1 Benchmark
We adopt the optimal offline solution, defined as the optimal matching cost without any future uncertainty, as our benchmark. For all set of servers $S$ and set of requests $R$ such that $|S|\geq|R|$, denote $\mathrm{OPT}(S,R)$ as the minimum cost of all matchings between $S$ and $R$ that match all requests. Given a set of servers $S$ with $|S|\geq n$ and a distribution for request arrival $\mathbb{D}$, the optimal offline solution on $S$ with $n$ requests randomly drawn from $\mathbb{D}$ is defined as $\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$, where $R\sim\mathbb{D}^{n}$ denotes that the requests in $R$ are independently drawn from $\mathbb{D}$. Furthermore, define $\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$.
We use competitive ratio and regret to measure the performance of online algorithms. For given metric space $(X,\delta)$ and distribution $\mathbb{D}$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances [^3] We require $|S|=n$ when considering balanced markets and $n\leq|S|\leq O(n)$ for unbalanced markets. $(S,n)$,
$\mathrm{cost}_{{\mathcal{A}}}(S,n)\leq\alpha\cdot\mathrm{OPT}(S,n)\,.$
Moreover, define the regret of an algorithm ${\mathcal{A}}$ for an instance $(S,n)$ as
$\mathrm{Regret}({\mathcal{A}};S,n)\triangleq\frac{1}{n}\left(\mathrm{cost}_{{\mathcal{A}}}(S,n)-\mathrm{OPT}(S,n)\right)\,.$
2.2 Optimal Offline Solutions for Euclidean Space
A considerable portion of this paper focuses on Euclidean space. In this subsection, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and bound the optimal offline solutions for smooth distributions and general distributions. All the omitted proofs in this subsection can be found in Appendix A.
For an arbitrary distribution $\mathbb{D}$, the following upper bounds for $\mathrm{OPT}(n)$ are known from the optimal transport literature.
Lemma 1 (Theorem 3.11 in [HSM22]).
For all $d\geq 1$ and distribution $\mathbb{D}$ on $[0,1]^{d}$,
$\mathrm{OPT}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n}\log n)&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
When restricted to smooth distributions, we improve the upper bound for $d=2$ in Lemma 1 by generalizing the proof of the upper bound in the Ajtai-Komlós-Tusnády Matching Theorem [AKT84], which asserts that $\mathrm{OPT}(n)=\Theta(\sqrt{n\log n})$ for uniform distribution $\mathbb{D}$ and $d=2$.
Lemma 2.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$,
$\mathrm{OPT}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Next, we establish the lower bounds for smooth distributions in balanced markets.
Lemma 3.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$,
$\mathrm{OPT}(n)\geq\begin{cases}\Omega(\sqrt{n})&d=1\\ \Omega(n^{1-\frac{1}{d}})&d\geq 2\end{cases}\,.$
The lower bound in Lemma 3 for $d=1$ follows from the quantile mapping and the well-known $\Omega(\sqrt{n})$ lower bound for the optimal offline solution with uniform distribution, and the lower bounds for $d\geq 2$ rely on nearest-neighbor-distance.
Finally, we present the lower bounds for smooth distributions in unbalanced markets, which are proved by employing nearest-neighbor-distance as well.
Lemma 4.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $n\leq|S|\leq O(n)$, $\mathrm{OPT}(S,n)\geq\Omega(n^{1-\frac{1}{d}})$.
2.3 The $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ Algorithm
The $\mathsf{Simulate\text{-}Optimize\text{-}Assign\text{-}Repeat}$ ($\mathsf{SOAR}$) algorithm is proposed by [CKKZ23] under the setting in which both servers and requests are stochastic. The $\mathsf{SOAR}$ algorithm, whose guarantee is stated below, will serve as an important building block of our algorithms and be invoked in a black-box manner.
Lemma 5 ([CKKZ23]).
Suppose that $m=n$ and $\mathbb{D}$ is a distribution on $X$, where the cost function $\delta$ is not required to satisfy the triangle inequality. When servers and requests are randomly drawn from $\mathbb{D}$, an algorithm exists with an expected cost of $\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)$.
3 Reducing Adversarial to Stochastic
In this section, we present an algorithmic framework to transform an algorithm for stochastic servers and requests into an algorithm for adversarial servers and stochastic requests at a moderate cost.
Theorem 1.
Given a distribution $\mathbb{D}$ on a metric space $(X,\delta)$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)$.
3.1 Proof of Theorem 1
Assuming that requests are stochastic, we start by lower bounding the optimal offline solution in balanced markets with adversarial servers. In particular, we show that it is at least the optimal offline solution in balanced markets with stochastic servers and requests up to a constant factor.
Lemma 6.
For all metric space $(X,\delta)$, for all distribution $\mathbb{D}$ over $X$ and set of servers $S^{\prime}$ with $|S^{\prime}|=n$, $\mathrm{OPT}(S^{\prime},n)\geq\frac{1}{2}\cdot\mathrm{OPT}(n)$.
Proof.
Given a set of servers $S^{\prime}$ with $|S^{\prime}|=n$, we claim that $\mathrm{OPT}(S,R)\leq\mathrm{OPT}(S,S^{\prime})+\mathrm{OPT}(S^{\prime},R)$ for all set of servers $S$ and set of requests $R$ with $|S|=|R|=n$. To see this, let $M_{1}$ be the perfect matching between $S$ and $S^{\prime}$ with the minimum cost and $M_{2}$ be the perfect matching between $S^{\prime}$ and $R$ with the minimum cost. Note that
$\mathrm{OPT}(S,S^{\prime})+\mathrm{OPT}(S^{\prime},R)$ $=\sum_{s^{\prime}\in S^{\prime}}\delta(M_{1}(s^{\prime}),s^{\prime})+\sum_{s^{\prime}\in S^{\prime}}\delta(s^{\prime},M_{2}(s^{\prime}))$
$\geq\sum_{s^{\prime}\in S^{\prime}}\delta(M_{1}(s^{\prime}),M_{2}(s^{\prime}))$
$\geq\mathrm{OPT}(S,R)\,,$
where the first inequality holds by the triangle inequality. Taking expectations on both sides over $S,R\sim\mathbb{D}^{n}$, we obtain
$\mathrm{OPT}(n)\leq\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,S^{\prime})+\mathrm{OPT}(S^{\prime},R)\right]=2\cdot\mathrm{OPT}(S^{\prime},n)\,,$
which concludes the proof. ∎
Now, we are ready to prove Theorem 1. Given a distribution $\mathbb{D}$ over a metric space $(X,\delta)$, let ${\mathcal{A}}$ be an algorithm such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$, and we construct an algorithm ${\mathcal{A}}^{\prime}$ for adversarial servers and stochastic requests as follows. Given a set of servers $S$, we randomly generate a set of points $I\sim\mathbb{D}^{n}$, and let $M^{S,I}$ be the matching between $S$ and $I$ with the minimum cost that matches all points in $I$. Then, we run ${\mathcal{A}}$ with the set of initial servers being $I$. For each new request $r$, suppose that ${\mathcal{A}}$ matches it to $i\in I$, and we match $r$ to $M^{S,I}(i)\in S$ in ${\mathcal{A}}^{\prime}$.
Next, we analyze the expected cost of ${\mathcal{A}}^{\prime}$. By the triangle inequality, the expected cost of ${\mathcal{A}}^{\prime}$ is upper bounded by
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathbb{E}_{I\sim\mathbb{D}^{n}}\left[\sum_{i\in I}\delta(M^{S,I}(i),i)+\mathrm{cost}_{{\mathcal{A}}}(I,n)\right]=\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)\,,$
which concludes the first part in Theorem 1. Furthermore, when $|S|=n$, by the guarantee of ${\mathcal{A}}$ and Lemma 6, we obtain
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\alpha\cdot\mathrm{OPT}(n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)\,,$
concluding the second part in Theorem 1.
4 Euclidean Space with Smooth Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and $\mathbb{D}$ to be $\beta$-smooth [^4] The smoothness of distributions is a mild technical assumption allowing us to lower bound $\mathrm{OPT}(n)$. To see the necessity, consider $\mathbb{D}$ a non-smooth distribution supported on a singleton, in which case $\mathrm{OPT}(n)=0$. for some constant $\beta>0$, and we will apply Theorem 1 to derive algorithms for both balanced and unbalanced markets. In Section 4.1, we focus on uniform distribution, which is a special case of smooth distributions, and give an $O(1)$-competitive algorithm for all $d\geq 1$ in balanced markets.
Next, we consider balanced markets with smooth distributions in Section 4.2. Recall that when $d=1$, an $O(1)$-competitive algorithm in balanced markets for general distributions [^5] In Section C, we also give an alternative $O(1)$-competitive algorithm for $d=1$ in balanced markets with general distributions. is given by [GGPW19]. Moreover, the above algorithm admits an incentive ratio of $O((\log\log\log n)^{2})$ for general metrics and distributions. However, for $d=2$ and smooth distributions, we are only able to estimate the optimal offline solution up to a $\Theta(\sqrt{\log n})$ factor, which prevents our approach from achieving an $o(\sqrt{\log n})$ competitive ratio. Hence, we only focus on the case where $d\geq 3$. In particular, we present an $O(1)$-competitive algorithm for all $d\geq 3$.
Finally, in Section 4.3, we extend our results for smooth distributions to unbalanced markets that possess at most linearly many excess servers, i.e., $n\leq m\leq O(n)$. We give an algorithm with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
4.1 Warm-up: Balanced Markets with Uniform Distribution
In this subsection, we consider balanced markets and assume $\mathbb{D}$ to be uniform distribution. We show that combining Theorem 1 with the Hierarchical Greedy algorithm of [Kan22] results in an $O(1)$-competitive algorithm for all $d\geq 1$.
We present our algorithm in the following theorem.
Theorem 2.
Suppose that $m=n$. For all $d\geq 1$, an $O(1)$-competitive algorithm exists for adversarial servers and uniformly distributed requests.
Proof.
Recall that [Kan22] gives an algorithm with a competitive ratio of $O(1)$ for all $d\geq 1$ when both servers and requests follow uniform distribution. Combining the above algorithm with Theorem 1 concludes the proof. ∎
4.2 Balanced Markets
In this subsection, we assume $d\geq 3$ and consider balanced markets with smooth distributions. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio for $d\geq 3$ when servers and requests are randomly drawn from $\mathbb{D}$. Together with Theorem 1, this leads to an $O(1)$-competitive algorithm for $d\geq 3$ with adversarial servers and stochastic requests.
We first present our main theorem in this subsection.
Theorem 3.
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are randomly drawn from $\mathbb{D}$, an $O(1)$-competitive algorithm exists for $d\geq 3$.
The rest of this section is devoted to proving Theorem 3. By Theorem 1, it suffices to show that the $\mathsf{SOAR}$ algorithm is $O(1)$-competitive for $d\geq 3$ when both servers and requests are randomly drawn from $\mathbb{D}$. We compute the expected cost of the $\mathsf{SOAR}$ algorithm in the following lemma. Even though we only consider $d\geq 3$ in this subsection, the costs of the $\mathsf{SOAR}$ algorithm for $d\leq 2$ computed by Lemma 7 will be useful afterward.
Lemma 7.
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. An algorithm ${\mathcal{A}}$ exists such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Proof.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5, and we estimate the cost of ${\mathcal{A}}$ for $d=1$, $d=2$, and $d\geq 3$ separately. Recall that an upper bound for $\mathrm{OPT}(n)$ is given by Lemma 2. For $d=1$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n}\right)\,.$
For $d=2$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\sqrt{\frac{\log t}{t}}\right)=O\left(\sqrt{\log n}\cdot\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n\log n}\right)\,.$
For $d\geq 3$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{t^{1/d}}\right)=O\left(\int_{1}^{n}\frac{1}{x^{1/d}}\mathrm{d}x\right)=O\left(n^{1-\frac{1}{d}}\right)\,.$
Combining all the above concludes the proof. ∎
Finally, combining Lemma 7 with the lower bounds for $\mathrm{OPT}(n)$ given by Lemma 3 concludes the proof of Theorem 3.
4.3 Unbalanced Markets
We extend our results for smooth distributions to unbalanced markets satisfying $n\leq m\leq O(n)$.
Theorem 4.
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are randomly drawn from $\mathbb{D}$, an algorithm exists with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
Proof.
By Lemma 7, an algorithm ${\mathcal{A}}$ exists such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Moreover, by Theorem 1, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. Combining the lower bounds for $\mathrm{OPT}(S,n)$ given by Lemma 4 concludes the proof. ∎
5 Euclidean Space with General Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance. We do not make any assumptions on the distribution and derive algorithms with nearly optimal regret bounds for all $d\geq 1$ in unbalanced markets with at most linearly many excess servers. In particular, our results recover the nearly optimal regret bounds by [CKKZ23], which is established in balanced markets with servers and requests following different distributions. In the weaker setting studied by [CKKZ23], the regret bound is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$, implying that our regret bounds are also nearly optimal.
Theorem 5.
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is an arbitrary distribution on $[0,1]^{d}$. An algorithm ${\mathcal{A}}^{\prime}$ exists that, for every set of servers $S$ with $|S|=m$, admits a regret of
$\mathrm{Regret}({\mathcal{A}}^{\prime};S,n)\leq\begin{cases}O(n^{-\frac{1}{2}})&d=1\\ O(n^{-\frac{1}{2}}\log n)&d=2\\ O(n^{-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Proof.
For an arbitrary distribution $\mathbb{D}$, by applying the same arguments as in the proof of Lemma 7 except that the upper bound for $\mathrm{OPT}(n)$ is given by Lemma 1 instead of Lemma 2, we obtain an algorithm ${\mathcal{A}}$ such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n}\log n)&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
By Theorem 1, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. As a result, the regret of ${\mathcal{A}}^{\prime}$ equals
$\mathrm{Regret}({\mathcal{A}}^{\prime};S,n)=\frac{1}{n}\left(\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)-\mathrm{OPT}(S,n)\right)\leq\frac{\mathrm{cost}_{{\mathcal{A}}}(n)}{n}\,,$
which concludes the proof. ∎
We briefly explain the reason behind the contrast that in unbalanced markets with $d\leq 2$, our approach admits a nearly optimal regret (Theorem 5) with a suboptimal competitive ratio (Theorem 4). In fact, our regret guarantee indicates that $\mathrm{cost}_{{\mathcal{A}}}(S,n)-\mathrm{OPT}(S,n)$ is small, yet we have no control over the lower bound for $\mathrm{OPT}(S,n)$ even with smooth distributions, which prevents us from achieving a nearly optimal competitive ratio.
6 Adversarial Setting
In this section, we focus on a more general setting where both servers and requests are adversarially chosen. This setting has been extensively studied in the past [Rag18, BBGN14], whereas all known guarantees are established in balanced markets. We show that the competitive ratio guarantees in balanced markets of algorithms satisfying certain natural conditions can be carried over to unbalanced markets, and the resulting competitive ratios in unbalanced markets only depend on the number of requests. Combined with the algorithms of prior work, we obtain competitive ratio guarantees in unbalanced markets with adversarial servers and requests, which in turn leads to improved guarantees in unbalanced markets with adversarial servers and stochastic requests. In particular, we show that both the $O(\log^{3}n)$-competitive algorithm of [MNP06] for general metrics and the $O(\log n)$-competitive algorithm of [GL12] for doubling metrics admit the same competitive ratios in unbalanced markets. The latter result also improves the $O(\sqrt{n})$ competitive ratio for $d=1$ in unbalanced markets with servers adversarially chosen and requests drawn from a smooth distribution in Theorem 4.
Given the execution of an online algorithm ${\mathcal{A}}$ on an instance with initial servers $S$, we use $N_{{\mathcal{A}}}(r\mid S)$ to denote the set of available servers closest to the request $r$ upon its arrival. We say that an online algorithm ${\mathcal{A}}$ is greedy if it always (randomly) matches a new request to one of the closest available servers, i.e., it matches $r$ to a server in $N_{{\mathcal{A}}}(r\mid S)$. In particular, every greedy algorithm is completely determined by a (randomized) tie-breaking rule such that when facing multiple closest available servers upon the arrival of a new request, the algorithm applies the tie-breaking rule to (randomly) decide who to match with. We say that the tie-breaking rule of a greedy algorithm ${\mathcal{A}}$ is local if it only depends on the set of closest available servers $N_{{\mathcal{A}}}(r\mid S)$. Moreover, the local tie-breaking rule of a greedy algorithm ${\mathcal{A}}$ is called non-increasing if the probability of selecting each server in $N_{{\mathcal{A}}}(r\mid S)$ does not increase by adding additional servers into $N_{{\mathcal{A}}}(r\mid S)$, i.e., for all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,.$
We show that the competitive ratio guarantee for every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets can be carried over to unbalanced markets.
Theorem 6.
For an arbitrary metric, if a greedy algorithm ${\mathcal{A}}$ with a local and non-increasing tie-breaking rule is $\alpha$-competitive in balanced markets with adversarial servers and requests, then ${\mathcal{A}}$ is $\alpha$-competitive in unbalanced markets with adversarial servers and requests.
We remark here that the resulting competitive ratio $\alpha$ for unbalanced markets can depend on the number of requests but is independent of the number of excess servers.
Proof of Theorem 6.
For two sets of points $S$ and $R$ in the metric with $|S|\geq\mid R|$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,R)$ as the expected cost of ${\mathcal{A}}$ when the set of initial servers is $S$ and the set of requests is $R$. Then, we claim that for all sets of points $S,S^{\prime},R$ in the metric with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, $\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$.
Before proving our claim, we first show the desired competitive ratio guarantee of ${\mathcal{A}}$ in unbalanced markets by applying the claim. Let $S$ be the set of initial servers and $R$ be the set of requests with $|S|\geq|R|=n$, and let $S^{\prime}\subseteq S$ satisfying $|S^{\prime}|=n$ and $\mathrm{OPT}(S,R)=\mathrm{OPT}(S^{\prime},R)$, i.e., $S^{\prime}$ consists of all matched servers in $S$ in the minimum cost matching between $S$ and $R$ that matches all requests in $R$. Since ${\mathcal{A}}$ is $\alpha$-competitive in balanced markets,
$\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S^{\prime},R)=\alpha\cdot\mathrm{OPT}(S,R)\,.$
Given that $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, by our claim,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S,R)\,,$
implying that ${\mathcal{A}}$ is $\alpha$-competitive in unbalanced markets.
It remains to prove our claim, which we establish by induction on $|R|$. When $|R|=0$, we have $\mathrm{cost}_{{\mathcal{A}}}(S,R)=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)=0$, and the statement straightforwardly holds. Assume for induction that the statement holds for $|R|=k$, and we show that the statement also holds for $|R|=k+1$. Given $S,S^{\prime},R$ with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|=k+1$, let $r\in R$ be the first arrived request. We use $M_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})$ to denote the (random) server that ${\mathcal{A}}$ matches $r$ to when the set of initial servers being $S$ and $S^{\prime}$, respectively. Since ${\mathcal{A}}$ is greedy, $M_{{\mathcal{A}}}(r\mid S)\in N_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. The following lemma states that either $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$ or $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$.
Lemma 8.
One of the following must hold: $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$ or $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$.
Proof.
Suppose for contradiction that $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})\neq\emptyset$ and $N_{{\mathcal{A}}}(r\mid S^{\prime})\not\subseteq N_{{\mathcal{A}}}(r\mid S)$, and let $s\in N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\setminus N_{{\mathcal{A}}}(r\mid S)$. Since $r$ is the first arrived request, all servers in $S$ (resp. $S^{\prime}$) are available upon the arrival of $r$, and hence, $N_{{\mathcal{A}}}(r\mid S)$ (resp. $N_{{\mathcal{A}}}(r\mid S^{\prime})$) contains all servers in $S$ (resp. $S^{\prime}$) closest to $r$. Moreover, note that $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq S^{\prime}\subseteq S$. On one hand, since $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, $\delta(s,r)=\delta(s^{\prime},r)$. On the other hand, since $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\notin N_{{\mathcal{A}}}(r\mid S)$, $\delta(s,r)<\delta(s^{\prime},r)$, leading to a contradiction. ∎
We analyze the expected cost of ${\mathcal{A}}$ under both cases in Lemma 8 separately. We start with the easier case where $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$, which implies that $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$ and $\delta(r,s)<\delta(r,s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. By $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$, we know that $M_{{\mathcal{A}}}(r\mid S)\notin S^{\prime}$ and hence $S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\}\subseteq S^{\prime}\subseteq S\setminus\{M_{{\mathcal{A}}}(r\mid S)\}$. By the inductive hypothesis, $\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})$. As a result,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)$ $=\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S))+\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})]$
$<\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S^{\prime}))+\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})]$
$=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\,,$
concluding the proof.
Next, we consider the case where $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$, which implies that $\delta(r,s)=\delta(r,s^{\prime})$ for all $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S)$. For every $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, define
$P_{S}(s)$ =$=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S)=s\right\},$ $Q_{S}(s)$ $=\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{s\},R\setminus\{r\})\,,$
$P_{S^{\prime}}(s^{\prime})$ =$=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S^{\prime})=s^{\prime}\right\},$ $Q_{S^{\prime}}(s^{\prime})$ $=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{s^{\prime}\},R\setminus\{r\})\,.$
Since the tie-breaking rule of ${\mathcal{A}}$ is local and non-increasing, $P_{S^{\prime}}(s)\geq P_{S}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. As a result,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)-\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$
$=\sum_{s\in N_{{\mathcal{A}}}(r\mid S)}P_{S}(s)\cdot(\delta(r,s)+Q_{S}(s))-\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S^{\prime}}(s)\cdot(\delta(r,s)+Q_{S^{\prime}}(s))$
$=\sum_{s\in N_{{\mathcal{A}}}(r\mid S)}P_{S}(s)\cdot Q_{S}(s)-\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S^{\prime}}(s)\cdot Q_{S^{\prime}}(s)$
$=\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}(P_{S}(s)\cdot Q_{S}(s)-P_{S^{\prime}}(s)\cdot Q_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)$
$\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s)\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)\,,$
where the second equality holds since $\delta(r,s)=\delta(r,s^{\prime})$ for all $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S)$ and the inequality holds since $Q_{S}(s)\leq Q_{S^{\prime}}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$ by the inductive hypothesis. Note that for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, we have $S^{\prime}\setminus\{s^{\prime}\}\subseteq S^{\prime}\subseteq S\setminus\{s\}$ and hence $Q_{S}(s)\leq Q_{S^{\prime}}(s^{\prime})$ by the inductive hypothesis. Let $B\triangleq\min_{s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s^{\prime})$, and it follows that $Q_{S}(s)\leq B\leq Q_{S^{\prime}}(s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. Therefore,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)-\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$ $\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s)\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)$
$\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}B\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot B$
$=B\cdot\left(\sum_{s\in N_{{\mathcal{A}}}(r\mid S)}P_{S}(s)-\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S^{\prime}}(s)\right)$
$=0\,,$
where the second inequality holds because $P_{S^{\prime}}(s)\geq P_{S}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. This concludes the proof. ∎
6.1 Applications on Hierarchically Separated Tree Metrics
Now, we apply Theorem 6 to algorithms of prior work. Recall that for balanced markets with adversarial servers and requests, the algorithm of [MNP06], abbreviated as the MNP algorithm, is $O(\log^{3}n)$-competitive for general metrics, and the Random-Subtree algorithm of [GL12] is $O(\log n)$-competitive for doubling metrics. We show that the competitive ratios of both algorithms are retained in unbalanced markets.
Both the MNP and Random Subtree algorithms are designed based on the same framework. They first randomly embed the original metric into a Hierarchically Separated Tree (HST) metric, which preserves the distance between every pair of nodes up to an $O(\log n)$ distortion. Next, they show that a randomized greedy subroutine with a specific tie-breaking rule admits a competitive ratio of $\beta$ for the HST metric. Overall, this leads to an $O(\beta\log n)$-competitive algorithm for the original metric. Hence, for our purpose, it suffices to show that the tie-breaking rule of the greedy subroutine for the HST metric is local and non-increasing.
Recall that an HST is a rooted tree satisfying the following property: For each subtree, the distances from the subtree’s root to all leaf nodes in the subtree are identical. Moreover, the metric embedding phase of the algorithms will map each location in the original metric to a leaf node of the HST. From now on, we will restrict our attention to the HST metric with all servers and requests locating at the HST’s leaf nodes.
We start by analyzing the greedy subroutine adopted by the MNP algorithm, which works as follows. Upon the arrival of a new request $r$, assuming the set of initial servers to be $S$, denote $A(r\mid S)$ as the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, and the subroutine matches $r$ to a uniformly random server among all available servers in the subtree rooted at $A(r\mid S)$. In the following lemma, we show that this tie-breaking rule is local and non-increasing, which implies that the greedy subroutine in the MNP algorithm satisfies the conditions in Theorem 6.
Lemma 9.
The greedy subroutine in the MNP algorithm admits a local and non-increasing tie-breaking rule.
Proof.
We use ${\mathcal{A}}$ to denote the greedy subroutine in the MNP algorithm. The tie-breaking rule of ${\mathcal{A}}$ is local by definition, and it remains to show that it is also non-increasing. Recall that $N_{{\mathcal{A}}}(r\mid S)$ is defined as the set of available servers closest to the request $r$ upon its arrival when the set of initial servers is $S$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}$ $=\frac{1}{|S^{\prime}|}\,,$
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}$ $=\frac{1}{|S^{\prime\prime}|}\,.$
Therefore,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,,$
which concludes that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing. ∎
The following corollary is a direct consequence of combining Theorem 6 and Lemma 9 with the competitive ratio guarantee of the MNP algorithm in balanced markets [MNP06].
Corollary 1.
Suppose that $m$ servers and $n$ requests are adversarially chosen with $n\leq m$. For an arbitrary metric, an $O(\log^{3}n)$-competitive algorithm exists.
We remark here that for general metrics with adversarial servers and requests, a randomized algorithm with an improved competitive ratio of $O(\log^{2}n)$ exists [BBGN14]. However, this algorithm does not belong to the family of greedy algorithms, and hence, Theorem 6 does not apply.
Next, we analyze the greedy subroutine adopted by the Random-Subtree algorithm, which works as follows. Upon the arrival of a new request $r$, recall that $A(r\mid S)$ denotes the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, assuming the set of initial servers to be $S$. Starting off at $A(r\mid S)$, the subroutine repeatedly moves us to a uniformly random subtree among all subtrees of the current node containing at least one available server until we reach a leaf node, and $r$ is matched to a uniformly random available server at the leaf node. In the following lemma, we show that this tie-breaking rule is local and non-increasing, which implies that the greedy subroutine in the Random Subtree algorithm satisfies the conditions in Theorem 6.
Lemma 10.
The greedy subroutine in the Random Subtree algorithm admits a local and non-increasing tie-breaking rule.
Proof.
We use ${\mathcal{A}}$ to denote the greedy subroutine in the Random Subtree algorithm. The tie-breaking rule of ${\mathcal{A}}$ is local by definition, and it remains to show that it is also non-increasing. For each node $x$ in the HST, denote $A_{\ell}(x)$ as the $\ell$-th ancestor of $x$ and $T_{\ell}(x)$ as the subtree rooted at $A_{\ell}(x)$. That is, $A_{0}(x)$ is $x$, $A_{1}(x)$ is the parent of $x$, and so on. Fix a request $r$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, let $\ell^{\prime}$ satisfy $A_{\ell^{\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$, and let $\ell^{\prime\prime}$ satisfy $A_{\ell^{\prime\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. Note that $\ell^{\prime}$ and $\ell^{\prime\prime}$ only depend on $S^{\prime}$ and $S^{\prime\prime}$, respectively.
We first show that $\ell^{\prime}=\ell^{\prime\prime}$ by deriving a contradiction when assuming $\ell^{\prime}\neq\ell^{\prime\prime}$. We only analyze the case where $\ell^{\prime}<\ell^{\prime\prime}$, and the case where $\ell^{\prime}>\ell^{\prime\prime}$ can be proved analogously. When $\ell^{\prime}<\ell^{\prime\prime}$, all servers in $S^{\prime}$ are located in $T_{\ell^{\prime}}(r)$, and all servers in $S^{\prime\prime}$ are located in $T_{\ell^{\prime\prime}}(r)$. Moreover, by the definition of $\ell^{\prime\prime}$, no server in $S^{\prime\prime}$ is located in $T_{\ell^{\prime\prime}-1}(r)$. Since $T_{\ell^{\prime}}(r)$ is a subtree in $T_{\ell^{\prime\prime}-1}(r)$, we have $S^{\prime}\cap S^{\prime\prime}=\emptyset$, contradicting the assumption that $S^{\prime}\subseteq S^{\prime\prime}$.
Fix a server $s\in S^{\prime}$. For every $i\in[\ell^{\prime}]$, denote $w_{i}^{\prime}$ as the number of subtrees of $A_{i}(s)$ containing at least one available server upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Moreover, define $w_{0}^{\prime}$ as the number of available servers collocated with $s$ (including $s$) upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Recall that $\ell^{\prime}=\ell^{\prime\prime}$, and define $w_{i}^{\prime\prime}$ for all $i\in\{0,1,\ldots,\ell^{\prime}\}$ similarly when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. By the definition of the tie-breaking rule of ${\mathcal{A}}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}$ $=\prod_{i=0}^{\ell^{\prime}}\frac{1}{w_{i}^{\prime}}\,,$
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}$ $=\prod_{i=0}^{\ell^{\prime}}\frac{1}{w_{i}^{\prime\prime}}\,.$
Since $S^{\prime}\subseteq S^{\prime\prime}$, we have $w_{i}^{\prime}\leq w_{i}^{\prime\prime}$ for all $i\in\{0,1,\ldots,\ell^{\prime}\}$. Therefore,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,,$
concluding that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing. ∎
Combining Theorem 6 and Lemma 10 with the competitive ratio guarantee of the Random Subtree algorithm in balanced markets [GL12], we derive the competitive ratio guarantee of the Random Subtree algorithm in unbalanced markets in the following corollary.
Corollary 2.
Suppose that $m$ servers and $n$ requests are adversarially chosen with $n\leq m$. For doubling metrics, an $O(\log n)$-competitive algorithm exists.
7 Discussion and Future Directions
In this paper, for balanced markets, we achieve an $O(1)$ competitive ratio for the Euclidean space with smooth distributions and $d\neq 2$. Nevertheless, it still remains open to give an $O(1)$-competitive algorithm for general metrics and distributions. The potential approaches for achieving this goal comprise (1) improving the analysis of the algorithm of [GGPW19], which is proved to be $O((\log\log\log n)^{2})$-competitive without any super-constant lower bound known, and (2) applying Theorem 1 and the $\mathsf{SOAR}$ algorithm to the simplified model given by [GGPW19] (Lemma 15). The main technical barrier underlying the latter approach lies in the difficulty of characterizing $\mathrm{OPT}(n)$ for general metrics and distributions.
We also derive an algorithm in unbalanced markets for the Euclidean space and smooth distributions with improved competitive ratios. Our algorithm admits an optimal constant competitive ratio for $d\geq 3$, leaving the improvement for $d\leq 2$ an intriguing future direction. Given that the line metric, i.e., $d=1$, is typically the most tractable case [GGPW19, BFP23], it is surprising that it becomes the most intractable one with the presence of excess servers, rendering its resolution more appealing.
Moreover, for adversarial servers and requests, we show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is preserved in unbalanced markets. However, our characterization fails to capture all the existing algorithms. In this most extensively studied setting, our understanding of the achievable guarantees in unbalanced markets remains very limited, and nothing beyond our results is known. An exciting future direction is to generalize the performance of the existing algorithms with provable guarantees in balanced markets to unbalanced markets.
Finally, we showcase the generality of our reduction by establishing positive results for cost functions only approximately satisfying the triangle inequality in Appendix B. Conversely, a strong impossibility result for achieving a bounded competitive ratio without any assumption on cost functions exists [GGPW19]. Our results can be viewed as pushing the boundary within which a bounded competitive ratio is achievable, and it would be interesting to further characterize the boundary between possibility and impossibility.
Acknowledgments
We thank Yash Kanoria, Akshit Kumar, Tristan Pollner, Shayan Talaei, and Yilun Chen for their helpful discussions, and Kiana Ashgri for pointing out the proof of Lemma 13.
Appendix A Postponed Proofs in Section 2.2
A.1 Proof of Lemma 2
The upper bounds for $d\neq 2$ directly follow from Lemma 1, and it remains to prove the upper bound for $d=2$. We use $\mathbb{U}$ to denote uniform distribution. Recall that the Ajtai-Komlós-Tusnády Matching Theorem [AKT84] states that $\mathbb{E}_{S,R\sim{\mathbb{U}}^{n}}[\mathrm{OPT}(S,R)]=\Theta(\sqrt{n\log n})$ for $d=2$. We will generalize the proof of the upper bound in the Ajtai-Komlós-Tusnády Matching Theorem interpreted by [Tal92, Section 4.3].
We first set up some notations necessary for the proof. For all metric space $(X,\delta)$ and $X^{\prime}\subseteq X$, let $\Delta(X^{\prime})$ denote the diameter of $X^{\prime}$ with respect to $\delta$, i.e.,
$\Delta(X^{\prime})=\sup_{x,y\in X^{\prime}}\delta\left(x,y\right)\,.$ (1)
Given a set $X$, an admissible sequence of $X$ is an increasing sequence $({\mathcal{A}}_{t})_{t\geq 0}$ of partitions of $X$ where the cardinality of ${\mathcal{A}}_{t}$ satisfies $|{\mathcal{A}}_{0}|=1$ and $|{\mathcal{A}}_{t}|\leq 2^{2^{t}}$ for $t\geq 1$. Given an admissible sequence $({\mathcal{A}}_{t})_{t\geq 0}$ of $X$, we use $A_{t}(x)$ to denote the unique element of ${\mathcal{A}}_{t}$ that contains $x$ for all $t\geq 0$ and $x\in X$. For all metric space $(X,\delta)$ and $\alpha>0$, define
$\gamma_{\alpha}\left(X,\delta\right)$ $\triangleq\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\Delta\left(A_{t}(x)\right)\,,$ (2)
where the infimum is taken over all admissible sequences of $X$.
Lemma 11.
For all $\alpha>0$, and metric spaces $(X,\delta_{1})$ and $(X,\delta_{2})$, if $\delta_{1}(x,y)\leq B\cdot\delta_{2}(x,y)$ for some $B>0$ and all $x,y\in X$, then $\gamma_{\alpha}\left(X,\delta_{1}\right)\leq B\cdot\gamma_{\alpha}\left(X,\delta_{2}\right)$.
Proof.
By (1) and (2),
$\gamma_{\alpha}\left(X,\delta_{1}\right)$ $=\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\left(\sup_{y,z\in A_{t}(x)}\delta_{1}\left(y,z\right)\right)$
$\leq B\cdot\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\left(\sup_{y,z\in A_{t}(x)}\delta_{2}\left(y,z\right)\right)=B\cdot\gamma_{\alpha}\left(X,\delta_{2}\right)\,,$
where the inequality holds by $\delta_{1}(x,y)\leq B\cdot\delta_{2}(x,y)$ for all $x,y\in X$. ∎
Now, we are ready to prove the upper bound for $d=2$. Let $X=[0,1]^{2}$, $\delta(x,y)=\left\|{x-y}\right\|_{2}$ for all $x,y\in X$, and $\mathbb{D}$ be a $\beta$-smooth distribution on $X$. Our goal is to prove that
$\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\leq O\left(\sqrt{n\log n}\right)\,.$
Let ${\mathcal{C}}$ denote the class of $1$-lipschitz functions on $(X,\delta)$, i.e., functions $f$ such that $|f(x)-f(y)|\leq\delta(x,y)$ for all $x,y\in X$. For all $f\in{\mathcal{C}}$ and $S=\{s_{1},\ldots,s_{n}\}\subseteq X$, define
$Z_{f}(S)\triangleq\sum_{i=1}^{n}\left(f(s_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)\,.$
By [Tal92, Proposition $4.2.3$], for all $S=\{s_{1},\ldots,s_{n}\},R=\{r_{1},\ldots,r_{n}\}\subseteq X$,
$\mathrm{OPT}(S,R)$ $=\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(s_{i})-f(r_{i})\right)$
$\leq\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(s_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)+\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(r_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)$
$=\sup_{f\in{\mathcal{C}}}|Z_{f}(S)|+\sup_{f\in{\mathcal{C}}}|Z_{f}(R)|\,.$ (3)
Analogous to [Tal92, Corollary $4.3.5$], for every $v>0$, we can show that
$\mathbb{P}_{S\sim\mathbb{D}^{n}}[\left|Z_{f}(S)\right|\geq v]\leq 2\exp\left(-\min\left(\frac{v^{2}}{4n\left\|{f}\right\|_{2}^{\mathbb{D}}},\frac{v}{4\left\|{f}\right\|_{\infty}^{\mathbb{D}}}\right)\right)\,,$ (4)
where $\left\|{f}\right\|_{p}^{\mathbb{D}}$ denotes the norm of $f$ in $L_{p}$ with respect to the measure $\mu_{\mathbb{D}}$ induced by the distribution $\mathbb{D}$. The proof of (4) is identical to the proof of [Tal92, Corollary $4.3.5$] by applying Bernstein’s inequality and hence omitted here. Then, analogous to [Tal92, Theorem $4.3.6$], for every class ${\mathcal{F}}$ of functions on $[0,1]^{2}$ with $0\in{\mathcal{F}}$, we can show that
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{F}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{F}},d_{2}^{\mathbb{D}}\right)+\gamma_{1}\left({\mathcal{F}},d_{\infty}^{\mathbb{D}}\right)\right)\,,$ (5)
where $d_{2}^{\mathbb{D}}$ and $d_{\infty}^{\mathbb{D}}$ are the distances induced on ${\mathcal{F}}$ by the norms of $L_{2}$ and $L_{\infty}$ respectively, with respect to the measure $\mu_{\mathbb{D}}$ induced by $\mathbb{D}$ on $[0,1]^{2}$, and $\gamma_{\alpha}$ is defined in (2) for $\alpha>0$. The proof of (5) is identical to the proof of [Tal92, Theorem $4.3.6$] by applying (4) and [Tal92, Theorem $2.2.23$], and hence omitted here.
Let $d_{2}^{\mathbb{U}}$ and $d_{\infty}^{\mathbb{U}}$ be the distances induced on ${\mathcal{F}}$ by the norms of $L_{2}$ and $L_{\infty}$ respectively, with respect to the measure induced by the uniform distribution on $[0,1]^{2}$. For all $f,g\in\mathcal{F}$,
$d_{2}^{\mathbb{D}}(f,g)$ $=\left(\int_{X}(f(x)-g(x))^{2}\mathrm{d}\mu_{\mathbb{D}}(x)\right)^{1/2}\leq\sqrt{\beta}\left(\int_{X}(f(x)-g(x))^{2}\mathrm{d}x\right)^{1/2}$
$=\sqrt{\beta}\left(\mathbb{E}_{x\sim\mathbb{U}}\left[(f(x)-g(x))^{2}\right]\right)^{1/2}=\sqrt{\beta}\cdot d_{2}^{\mathbb{U}}(f,g)\,,$
where the inequality holds by the assumption that $\mathbb{D}$ is $\beta$-smooth. Also, note that
$d_{\infty}^{\mathbb{D}}(f,g)\leq d_{\infty}^{\mathbb{U}}(f,g)=\sup_{x\in X}|f(x)-g(x)|\,.$
Then, by Lemma 11 and (5), together with the above two displayed inequalities, given that $\beta>0$ is a constant, we have
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{F}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{F}},d_{2}^{\mathbb{U}}\right)+\gamma_{1}\left({\mathcal{F}},d_{\infty}^{\mathbb{U}}\right)\right)\,.$
Recall that ${\mathcal{C}}$ denotes the class of $1$-lipschitz functions on $(X,\delta)$. Then, we obtain
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{C}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{C}},d_{2}^{\mathbb{U}}\right)+\gamma_{1}\left({\mathcal{C}},d_{\infty}^{\mathbb{U}}\right)\right)\leq O\left(\sqrt{n\log n}\right)\,,$ (6)
where the second inequality is implicitly shown in the proof of [Tal92, Theorem $4.3.2$]. Combining (3) and (6), it holds that
$\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\leq 2\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{C}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{n\log n}\right)\,,$
which concludes the proof.
A.2 Proof of Lemma 3
For $d=1$, let $F(x)=\int_{0}^{x}\mu_{\mathbb{D}}(x)\mathrm{d}x$ for $x\in[0,1]$ be the cumulative density function of $\mathbb{D}$. Let $S=\{s_{1},\ldots,s_{n}\}\sim\mathbb{D}^{n}$ and $R=\{r_{1},\ldots,r_{n}\}\sim\mathbb{D}^{n}$ be two sets of random points drawn from $\mathbb{D}$. Define $P_{S}=\{F(s_{1}),\ldots,F(s_{n})\}$ and $P_{R}=\{F(r_{1}),\ldots,F(r_{n})\}$, and each point in $P_{S}\cup P_{R}$ follows uniform distribution on $[0,1]$. Since the expected cost of the optimal matching between two sets of $n$ uniformly distributed points on $[0,1]$ is $\Theta(\sqrt{n})$, we have
$\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]=\Theta\left(\sqrt{n}\right)\,.$ (7)
Note that for all $s,r\in[0,1]$,
$|F(s)-F(r)|=\left|\int_{r}^{s}f(x)\mathrm{d}x\right|\leq\left|\int_{r}^{s}\beta\mathrm{d}x\right|=\beta\cdot|s-r|\,.$ (8)
Let $M$ be the perfect matching between $S$ and $R$ with the minimum cost, and let $M^{\prime}$ be the perfect matching between $P_{S}$ and $P_{R}$ induced by $M$ such that $F(s)\in P_{S}$ is matched with $F(M(s))\in P_{R}$ for each $s\in S$. Hence, the cost of $M^{\prime}$ equals
$\sum_{s\in S}|F(s)-F(M(s))|\geq\mathrm{OPT}(P_{S},P_{R})\,.$ (9)
As a result,
$\mathrm{OPT}(S,R)=\sum_{s\in S}|s-M(s)|\geq\sum_{s\in S}\frac{1}{\beta}\cdot|F(s)-F(M(s))|\geq\frac{1}{\beta}\cdot\mathrm{OPT}(P_{S},P_{R})\,,$
where the equality holds by the definition of $M$, the first inequality holds by (8), and the second inequality holds by (9). Taking expectations on both sides, we obtain
$\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq\frac{1}{\beta}\cdot\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]\geq\Omega\left(\sqrt{n}\right)\,,$
where the last inequality holds by (7). This concludes the lower bound for $d=1$.
The lower bounds for $d\geq 2$ are proved by employing nearest-neighbor-distance, which is provided in the following lemma.
Lemma 12.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $|S|=n$,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(n^{-1/d}\right)\,.$
Proof.
Fix $S\subseteq[0,1]^{d}$ with $|S|=n$. For all $s\in S$ and $\ell>0$, let $B(s,\ell)=\{x\in[0,1]^{d}\mid\left\|{x-s}\right\|_{2}\leq\ell\}$ denote the ball restricted in $[0,1]^{d}$ with center $s$ and radius $\ell$, and we have
$\mathbb{P}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\leq\ell]=\int_{B(s,\ell)}\mu_{\mathbb{D}}(x)\mathrm{d}x\leq\int_{B(s,\ell)}\beta\mathrm{d}x\leq C\cdot\ell^{d}$
for some constant $C>0$. For all $s\in S$ and $r\in[0,1]^{d}$, let $O_{s,r}$ denote the event that $s\in\mathop{\arg\min}_{s^{\prime}\in S}\left\|{s^{\prime}-r}\right\|_{2}$ and $p_{s}=\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]$. For all $s\in S$ and $0<\ell\leq(p_{s}/C)^{1/d}$, it holds that
$\mathbb{P}_{r\sim\mathbb{D}}\left[O_{s,r}\land(\left\|{s-r}\right\|_{2}\geq\ell)\right]$ $=\mathbb{P}_{r\sim\mathbb{D}}\left[O_{s,r}\right]-\mathbb{P}_{r\sim\mathbb{D}}\left[O_{s,r}\land(\left\|{s-r}\right\|_{2}\leq\ell)\right]$
$\geq p_{s}-\mathbb{P}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\leq\ell\right]$
$\geq p_{s}-C\cdot\ell^{d}\,.$
As a result, for every $s\in S$,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\mid O_{s,r}\right]$ $\geq\int_{0}^{(p_{s}/C)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\geq\ell\mid O_{s,r}\right]\mathrm{d}\ell$
$=\frac{1}{p_{s}}\int_{0}^{(p_{s}/C)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}\land(\left\|{s-r}\right\|_{2}\geq\ell)]\mathrm{d}\ell$
$\geq\frac{1}{p_{s}}\int_{0}^{(p_{s}/C)^{1/d}}(p_{s}-C\cdot\ell^{d})\mathrm{d}\ell$
$=\frac{d}{d+1}\cdot\left(\frac{p_{s}}{C}\right)^{1/d}\,.$
Therefore,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]=\sum_{s\in S}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]\cdot\mathbb{E}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\mid O_{s,r}]\geq\sum_{s\in S}\Omega\left(p_{s}^{1+\frac{1}{d}}\right)\geq\Omega\left(n^{-1/d}\right)\,,$
where the last inequality holds since $\sum_{s\in S}p_{s}=1$ and $p_{s}\geq 0$ for every $s\in S$. ∎
Now, we are ready to prove the lower bounds for $d\geq 2$. By Lemma 12,
$\mathrm{OPT}(n)$ $=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]=\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\right]$
$\geq\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\right]\geq\Omega\left(n^{1-\frac{1}{d}}\right)\,,$
which concludes the proof.
A.3 Proof of Lemma 4
By Lemma 12 and the assumption that $|S|\leq O(n)$,
$\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(n^{1-\frac{1}{d}}\right)$
as desired.
Appendix B Cost Functions Beyond Triangle Inequality
In previous sections, we consider metric space where the cost function satisfies the exact triangle inequality. In this section, we showcase the generality of our reduction in Section 3 by proving positive results when the cost function only approximately satisfies the triangle inequality. Specifically, we say that a cost function $\delta$ on $X$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z\in X$,
$\delta(x,z)\leq\eta\cdot(\delta(x,y)+\delta(y,z))\,.$
We say that a metric space satisfies the $\eta$-approximate triangle inequality if the corresponding cost function satisfies the $\eta$-approximate triangle inequality.
We first generalize Theorem 1 to the case where the metric space only satisfies the approximate triangle inequality.
Theorem 7.
Given a distribution $\mathbb{D}$ on a metric space $(X,\delta)$ satisfying the $\eta$-approximate triangle inequality for $\eta>0$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\eta\cdot(\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n))$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha\eta^{2}+\eta)\cdot\mathrm{OPT}(S,n)$.
The proof of Theorem 7 can be obtained by replacing every usage of the triangle inequality by the approximate triangle inequality in the proof of Theorem 1 and hence is omitted.
Next, we consider the case where $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\geq 1$, which we will prove that satisfies the $2^{p-1}$-approximate triangle inequality [^6] Here, $p$ can be a real number., and assume $\mathbb{D}$ to be uniform distribution. We show that in balanced markets with adversarial servers and stochastic requests, an $O(1)$-competitive algorithm exists for $d\geq 1$ and $p\in[1,\max\{d,2\})$.
Theorem 8.
Assume that $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $d\geq 1$ and $p\in[1,\max\{d,2\})$. Supposing that $m=n$ and $\mathbb{D}$ is uniform distribution, an $O(1)$-competitive algorithm exists with adversarial servers and stochastic requests.
We remark here that competitive ratio guarantees for other combinations of $d$ and $p$, more general distributions, and unbalanced markets can be similarly achieved as in Section 4 and Section 5. Nevertheless, we didn’t pursue this direction since parts of the resulting competitive ratios seem suboptimal, i.e., super-constant, and no new algorithmic insight emerges.
B.1 Proof of Theorem 8
We start by showing that for all $p\geq 1$, the cost function $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ satisfies the $2^{p-1}$-approximate triangle inequality by proving a more general statement: Assuming $\delta^{\prime}$ to be a cost function satisfying the triangle inequality, the cost function defined as $\delta(x,y)\triangleq\delta^{\prime}(x,y)^{p}$ satisfies the $2^{p-1}$-approximate triangle inequality.
Lemma 13.
For every $p\geq 1$, given a cost function $\delta^{\prime}$ on $X$ satisfying the triangle inequality, the cost function on $X$ defined as $\delta(x,y)\triangleq\delta^{\prime}(x,y)^{p}$ for all $x,y\in X$ satisfies the $2^{p-1}$-approximate triangle inequality.
Proof.
For all $x,y,z\in X$,
$\delta(x,z)$ $=\delta^{\prime}(x,z)^{p}\leq(\delta^{\prime}(x,y)+\delta^{\prime}(y,z))^{p}=\left(\frac{\delta^{\prime}(x,y)}{2}+\frac{\delta^{\prime}(y,z)}{2}\right)^{p}\cdot 2^{p}$
$\leq\left(\frac{\delta^{\prime}(x,y)^{p}}{2}+\frac{\delta^{\prime}(y,z)^{p}}{2}\right)\cdot 2^{p}=(\delta(x,y)+\delta(y,z))\cdot 2^{p-1}\,,$
where the first inequality holds by the triangle inequality and the second inequality holds by the convexity of $g(x)=x^{p}$ with $p\geq 1$. ∎
Next, we present the following estimation of $\mathrm{OPT}(n)$.
Lemma 14 ([Led19]).
Let $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $d,p\geq 1$. Assuming ${\mathbb{D}}$ to be uniform distribution, it holds that
$\mathrm{OPT}(n)=\begin{cases}\Theta(n^{1-\frac{p}{2}})&d=1\\ \Theta(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}})&d=2\\ \Theta(n^{1-\frac{p}{d}})&d\geq 3\end{cases}\,.$
Now, we are ready to prove Theorem 8. Let ${\mathcal{A}}$ be the algorithm as per Lemma 5, and we analyze $\mathrm{cost}_{{\mathcal{A}}}(n)$ for $d=1$, $d=2$, and $d\geq 3$ separately by applying Lemma 14. For $d=1$ and $p\in[1,2)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{2}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
For $d=2$ and $p\in[1,2)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}(\log t)^{\frac{p}{2}}\right)=O\left((\log n)^{\frac{p}{2}}\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)$
$=O\left(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
For $d\geq 3$ and $p\in[1,d)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{d}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{d}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{d}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
As a result, ${\mathcal{A}}$ admits an $O(1)$ competitive ratio when both servers and requests are uniformly distributed. By Lemma 13, $\delta$ satisfies the $2^{p-1}$-approximate triangle inequality. Combined with Theorem 7, we conclude that an $O(1)$-competitive algorithm exists in balanced markets with adversarial servers and uniformly distributed requests.
Appendix C An Alternative Algorithm for Tree Metrics
In this section, we consider balanced markets and apply our approach to tree metrics with general distributions. In particular, a tree metric is defined by shortest path distances in a tree $T=(V,E)$. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio with both servers and requests randomly drawn from $\mathbb{D}$, which implies the existence of an $O(1)$-competitive algorithm with servers adversarially chosen and requests randomly drawn from $\mathbb{D}$ by Theorem 1. This gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions, recovering the result by [GGPW19].
We first present our main theorem in this section.
Theorem 9.
Suppose that $m=n$ and consider tree metrics. For an arbitrary distribution $\mathbb{D}$, an $O(1)$-competitive algorithm exists when servers are adversarially chosen and requests are randomly drawn from $\mathbb{D}$.
To prove Theorem 9, it is known that in balanced markets with adversarial servers and stochastic requests following an arbitrary distribution, only a constant factor in the competitive ratio is lost by assuming that each request is uniformly distributed over server locations. That is, for every server $s\in S$, the request is collocated with $s$ with probability $1/n$.
Lemma 15 ([GGPW19]).
Suppose that $m=n$ and servers are adversarially chosen. For an arbitrary metric, if an $\alpha$-competitive algorithm exists when requests are uniformly distributed over server locations, then we can construct a $(2\alpha+1)$-competitive algorithm when requests follow an arbitrary distribution.
Given Lemma 15, it suffices to consider distribution $\mathbb{D}$ of the following form: Given a tree metric $T=(V,E)$, each node $v\in V$ is associated with an integer $a_{v}\geq 0$ such that $\sum_{v\in V}a_{v}=n$ and $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$. For each edge $e\in E$, let $d_{e}$ denote the length of $e$, and let $T_{1}(e)$ and $T_{2}(e)$ be two subtrees of $T$ obtained by deleting $e$. For each $e\in E$, define $n_{e}=\sum_{v\in T_{1}(e)}a_{v}$. Given such a distribution $\mathbb{D}$, we estimate $\mathrm{OPT}(t)$ in the following lemma.
Lemma 16.
For each $t\in[n]$, $\mathrm{OPT}(t)=\Theta\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)$.
Proof.
Fix $t\in[n]$. For each $e\in E$, let $P_{e},Q_{e}\sim{\rm Binom}(t,n_{e}/n)$ be random variables denoting the number of servers located in $T_{1}(e)$ and the number of requests located in $T_{1}(e)$, respectively. Given a realization of $(P_{e},Q_{e})_{e\in E}$, in the perfect matching between servers and requests with the minimum cost, for each $e\in E$, exactly $|P_{e}-Q_{e}|$ matched server-request pairs will utilize $e$, incurring a cost of $d_{e}\cdot|P_{e}-Q_{e}|$. Hence,
$\mathrm{OPT}(t)=\mathbb{E}\left[\sum_{e\in E}d_{e}\cdot|P_{e}-Q_{e}|\right]=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\,.$
We first lower bound $\mathrm{OPT}(t)$, which relies on the following probabilistic bound.
Claim 1 ([BK13]).
Let $Z\sim{\rm Binom}(n,p)$, with $n\geq 2$ and $p\in[1/n,1-1/n]$. Then, we have
$\mathbb{E}[|Z-\mathbb{E}Z|]\geq\mathrm{std}(Z)/\sqrt{2}\,,$
where $\mathrm{std}(Z)$ denotes the standard deviation of $Z$.
We lower bound $\mathrm{OPT}(t)$ by
$\mathrm{OPT}(t)$ $=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\geq\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-\mathbb{E}P_{e}|]$
$\geq\frac{1}{\sqrt{2}}\sum_{e\in E}d_{e}\cdot\mathrm{std}(P_{e})=\Omega\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)\,,$
where the first inequality holds since the function $f(z)=\mathbb{E}[|P_{e}-z|]$ reaches its minimum when $z=\mathbb{E}P_{e}$, and the second inequality holds by Claim 1. Moreover, $\mathrm{OPT}(t)$ is upper bounded by
$\mathrm{OPT}(t)$ $=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\leq\sum_{e\in E}d_{e}\left(\mathbb{E}\left[\left|P_{e}-\mathbb{E}P_{e}\right|\right]+\mathbb{E}\left[\left|Q_{e}-\mathbb{E}Q_{e}\right|\right]\right)$
$\leq\sum_{e\in E}d_{e}(\mathrm{std}(P_{e})+\mathrm{std}(Q_{e}))=O\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)\,,$
where the first inequality holds by $\mathbb{E}P_{e}=\mathbb{E}Q_{e}$ for every $e\in E$. ∎
Now, we are ready to prove Theorem 9.
Proof of Theorem 9.
According to Lemma 15, we can assume requests to be uniformly distributed over server locations by losing a constant factor in the competitive ratio. In other words, let $a_{v}$ be the number of servers located at $v\in V$, the distribution $\mathbb{D}$ satisfies $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$ for every $v\in V$. By Theorem 1, it suffices to give an $O(1)$-competitive algorithm when both servers and requests are randomly drawn from $\mathbb{D}$.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5. Recall that $\mathrm{OPT}(t)$ is estimated by Lemma 16. When both servers and requests are randomly drawn from $\mathbb{D}$, the expected cost of ${\mathcal{A}}$ equals
$\mathrm{cost}_{{\mathcal{A}}}(n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{n\sqrt{t}}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)$
$=O\left(\frac{1}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\frac{1}{\sqrt{n}}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)=O\left(\mathrm{OPT}(n)\right)\,,$
concluding that ${\mathcal{A}}$ is $O(1)$-competitive. ∎