### v2

Abstract
We study the online metric matching problem. There are $m$ servers and $n$ requests located in a metric space, where all servers are available upfront and requests arrive one at a time. Upon the arrival of a new request, it needs to be immediately and irrevocably matched to a free server, resulting in a cost of their distance. The objective is to minimize the total matching cost.
When servers are adversarial and requests are independently drawn from a known distribution, we reduce the problem to a more accessible setting where all servers and requests are independently drawn from the same distribution. Applying our reduction, for the Euclidean space $[0,1]^{d}$ with various choices of distributions, we achieve improved competitive ratios and nearly optimal regrets in both balanced and unbalanced markets. In particular, we give $O(1)$-competitive algorithms for $d\geq 3$ in both balanced and unbalanced markets with smooth distributions. Our algorithms improve on the $O((\log\log\log n)^{2})$ competitive ratio of Gupta et al. (ICALP’19) for balanced markets in various regimes, and provide the first positive results for unbalanced markets.
Moreover, when the servers and requests are adversarial and a prediction of request locations is provided, for balanced markets, we present a general framework for transforming an arbitrary algorithm that does not use predictions into an algorithm that leverages predictions. The competitive ratio of the resulting algorithm is a constant with a perfect prediction and degrades smoothly as the prediction accuracy deteriorates, while preserving the worst-case guarantee.
1 Introduction
The development of centralized matching markets has emerged as a crucial research area across operations research, economics, and computer science, driven by its diverse applications in kidney exchange programs, ride-sharing services, and online advertising platforms. Market designers in these domains face the fundamental challenge of making matching decisions under future uncertainty, with the primary objective of maximizing the total incurred value. The growing significance of these applications has sparked extensive research across multiple related areas. These include online bipartite matching [KVV90, Meh13, HTW24], online resource allocation [VB21, BKK23, BLM23, BBP24, Bra24, JMZ24], and dynamic matching [AAGK17, ALG20, WXY23, Gup24, KAG24, KYY25].
A particularly compelling example of these applications can be found in the recent proliferation of ride-hailing platforms such as Didi, Uber, and Lyft. In these systems, the platform must make an immediate matching decision when a customer requests a ride - significantly delaying or rejecting the request is typically not an option. The primary cost metric in each match is the distance a driver needs to travel to reach their assigned customer. While a simple greedy approach matches each arriving customer to their nearest available driver, such myopic decisions can often lead to suboptimal outcomes. For instance, after matching a driver to their nearest customer, the next customer request might arrive in that driver’s original location, requiring a much longer pickup distance from the remaining available drivers. Therefore, the platform’s key challenge is to design matching policies that minimize the total pickup cost by considering both current matches and potential future request patterns, while maintaining the constraint of immediate matching for each request.
Inspired by the special features of the market dynamics underlying ride-hailing platforms, in this paper, we focus on the online (bipartite) minimum cost matching problem. In this problem, $m$ nodes from one side of the market are called servers and are fully known upfront. On the other side of the market, which is often referred to as the request side, $n$ nodes arrive sequentially. We model the instantaneity of performing a match by assuming that each arriving request has to be immediately and irrevocably assigned to an available server. Due to the importance of application to ride-hailing, we assume that all servers and requests are located in the same metric space, and the cost incurred by matching a server and a request is measured by the distance between them. After realizing a match, both the matched server and request are removed. The objective of a central market planer is to minimize the total matching cost. We will consider both balanced and unbalanced markets, where balanced markets contain an equal number of servers and requests, and unbalanced markets possess at most a linear amount of excess servers. We will adopt competitive ratio (resp. regret) as our benchmark, which is defined as the ratio (resp. difference) between the expected average cost of matching each request incurred by the algorithm and that incurred by the optimal offline solution.
A majority of literature on online metric matching focuses on the adversarial model [PS23, Rag18, MNP06], where the locations of all servers and requests are adversarially chosen, or the weaker random-order model [Rag16], which assumes that the arrival order of adversarially chosen requests is randomly permuted. Both models, albeit being fairly general, suffer from super-constant competitive ratio lower bounds [Rag16, PS23]. Recent lines of work bypass these barriers by modeling more practical and tractable scenarios in which algorithms are equipped with additional information regarding future uncertainty. Specifically, [GGPW19, BFP23] assume that requests are randomly drawn from a known distribution, and [ACE+23b] consider the case where algorithms are provided with predictions of request locations. They achieve improved or even nearly optimal results under various choices of modeling assumptions.
1.1 Main Contribution
As our main contribution, we propose a generic approach to designing algorithms for the online metric matching problem that leverages information about future uncertainty - either through probability distributions or predictions of request locations. These algorithms result from enhancing an arbitrary algorithm ${\mathcal{A}}$ that does not use additional information. A key advantage of our approach is that this enhancement only invokes ${\mathcal{A}}$ in a black-box manner without imposing technical assumptions on ${\mathcal{A}}$.
More specifically, given predictions of request locations, for an arbitrary algorithm ${\mathcal{A}}$ that does not use predictions, we can convert ${\mathcal{A}}$ into an enhanced algorithm ${\mathcal{A}}^{\prime}$ that bases its matching decisions on both the predictions and ${\mathcal{A}}$’s strategy. The performance of ${\mathcal{A}}^{\prime}$ depends jointly on the prediction accuracy and ${\mathcal{A}}$’s performance (Theorem 1). When given distributional information instead, we generate predictions by sampling from the distribution, allowing us to apply the same approach.
We then demonstrate the effectiveness of our framework through several significant results in the stochastic and prediction settings respectively, following by our additional results for the adversarial setting.
Stochastic Setting.
When servers are adversarially chosen and requests are independently drawn from a known distribution $\mathbb{D}$, we show that the aforementioned approach results in a reduction to the setting with stochastic servers and requests. Specifically, if there is an algorithm ${\mathcal{A}}$ for balanced markets with servers and requests all independently drawn from $\mathbb{D}$, we can use ${\mathcal{A}}$ in a black-box manner to derive another algorithm ${\mathcal{A}}^{\prime}$ for unbalanced markets with servers adversarially chosen and requests independently drawn from $\mathbb{D}$ (Theorem 2). Moreover, if the markets are balanced, which can be viewed as a special case of unbalanced markets, ${\mathcal{A}}^{\prime}$ admits the same competitive ratio as ${\mathcal{A}}$ up to a constant factor. Next, we describe our results in more concrete settings obtained by further utilizing the recent progress in the setting with stochastic servers and requests [Kan22, CKKZ23].
Balanced Markets Unbalanced Markets
Uniform Dis. Smooth Dis. General Dis. Smooth Dis. General Dis.
$d=1$ $O(1)$ $O(1)$ $O(1)$ $O(\log n)$ $O(\log n)$
$d=2$ $O(1)$ ($\omega(1)$) $\omega(1)$ $\omega(1)$ $O(\sqrt{\log n})$ $O(\log n)$
$d\geq 3$ $O(1)$ ($\omega(1)$) $O(1)$ ($\omega(1)$) $\omega(1)$ $O(1)$ $O(\log n)$
Table 1: Competitive ratios for the Euclidean space $[0,1]^{d}$ with adversarial servers and stochastic requests. All previous state-of-the-art results are given in black fonts, where we use $\omega(1)$ to denote $O((\log\log\log n)^{2})$, and all new results achieved in this paper are marked in red. In particular, for balanced markets with general distributions, the competitive ratios of $O(1)$ for $d=1$ and of $O((\log\log\log n)^{2})$ for $d\geq 2$ are obtained by [GGPW19], which imply the bounds on their left.
We first introduce the competitive ratio results for the Euclidean space $[0,1]^{d}$ achieved by combining our reduction with algorithms of prior work, which are summarized in Table 1. Prior to our work, the best algorithm in balanced markets with general metrics and distributions achieves a competitive ratio of $O((\log\log\log n)^{2})$, and the same algorithm is known to be $O(1)$-competitive for tree metrics [GGPW19]. We start by showing that combining our reduction with the Hierarchical Greedy algorithm of [Kan22] leads to an $O(1)$-competitive algorithm in balanced markets for all $d\geq 1$ with adversarial servers and uniformly distributed requests (Theorem 4).
Next, we describe our results in balanced markets with servers adversarially chosen and requests independently drawn from a smooth distribution, where a distribution is smooth if it admits a density function upper bounded by a constant. For the algorithm ${\mathcal{A}}$ required by the reduction, we adopt the Simulate-Optimize-Assign-Repeat (SOAR) algorithm of [CKKZ23], which results in an $O(1)$-competitive algorithm for $d\geq 3$ (Theorem 5).
Finally, we present our results in unbalanced markets with adversarial servers and stochastic requests. When requests are independently drawn from a smooth distribution, by applying the same approach as for balanced markets, the resulting algorithm admits a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$ (Theorem 6). To improve the undesirable competitive ratio for $d=1$, we show that the $O(\log n)$ competitive ratio of the algorithm proposed by [GL12], which holds in balanced markets when all servers and requests are adversarially located in a doubling metric [^1] A metric space is said to be doubling if a doubling constant $M>0$ exists such that for all $r>0$, every ball of radius $r$ can be covered by at most $M$ balls of radius $r/2$. In particular, the Euclidean space $[0,1]^{d}$ is a doubling metric for constant $d$., is preserved in unbalanced markets (Corollary 3). These constitute the first known results in unbalanced markets for adversarial servers and stochastic requests.
We further describe our additional results with respect to regret analysis, which are for the Euclidean space $[0,1]^{d}$. Recall that in balanced markets with servers and requests following different and arbitrary distributions, [CKKZ23] show that the SOAR algorithm admits a regret of $O(n^{-\frac{1}{2}})$ for $d=1$, $O(n^{-\frac{1}{2}}\log n)$ for $d=2$, and $O(n^{-\frac{1}{d}})$ for $d\geq 3$, which is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$. By combining our reduction with the SOAR algorithm, we show that the same regrets are achievable even in unbalanced markets with servers adversarially chosen and requests independently drawn from an arbitrary distribution (Theorem 7). Since balanced markets form a special case of unbalanced markets, and our setting with adversarial servers is more general, the lower bounds by [CKKZ23] continue to hold in our case, implying our regret guarantees to be nearly optimal.
Prediction Setting.
In the prediction setting, algorithms receive a set of $n$ points as a prediction at the beginning, which predicts the unknown request locations. We will focus on balanced markets with servers and requests being adversarially chosen. In particular, the prediction is not guaranteed to be correct with its error oblivious to algorithms, and algorithms are expected to be (1) consistent: the performance of algorithms is nearly optimal with a perfect prediction, and (2) robust: when the prediction admits a large error, the performance of algorithms is comparable to that of the best online algorithm that does not use predictions. In our case, the prediction error $\eta$ is defined as the minimum cost of all perfect matchings between the set of predicted request locations and the set of actual request locations.
Given an $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions, we apply our generic approach mentioned at the beginning of this subsection to derive an algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions and achieves a competitive ratio of $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$, where $\mathrm{OPT}$ denotes the optimal offline cost (Theorem 8). To see the consistency of ${\mathcal{A}}^{\prime}$, its competitive ratio converges to a constant when the prediction error $\eta$ approaches 00. Moreover, the competitive ratio of ${\mathcal{A}}^{\prime}$ degrades smoothly as the prediction error increases, and the competitive ratio guarantee of $\alpha$ provided by ${\mathcal{A}}$ is maintained up to a constant factor regardless of the prediction error, indicating the robustness of ${\mathcal{A}}^{\prime}$. Finally, we apply the above result to several specific settings (Corollary 1).
Adversarial Setting.
Now, we discuss our results for the setting in which servers and requests are adversarially chosen with no predictions provided. As alluded to previously, we generalize the $O(\log n)$ competitive ratio of the algorithm by [GL12] for doubling metrics from balanced markets to unbalanced markets. This is achieved based on the observation that every greedy algorithm with a tie-breaking rule satisfying certain properties admits the same competitive ratio in balanced and unbalanced markets (Theorem 9). As a by-product, we similarly show that the $O(\log^{3}n)$ competitive ratio of the algorithm by [MNP06] in balanced markets with servers and requests adversarially located in an arbitrary metric is retained in unbalanced markets (Corollary 2). We remark here that all prior results for adversarial servers and requests are established in balanced markets, and it is not immediate whether they can be carried over to unbalanced markets.
1.2 Related Work
Online Metric Matching.
A majority of literature for online metric matching focuses on the adversarial setting in balanced markets. For general metrics with adversarial servers and requests, [KP93], [KMV94], and [Rag16] respectively give $(2n-1)$-competitive deterministic algorithms, which is also shown to be tight. When randomization is allowed, an $O(\log^{3}n)$-competitive algorithm is proposed by [MNP06]; it is subsequently improved by [BBGN14] to $O(\log^{2}n)$, which remains the best upper bound. In the random-order model, the deterministic algorithm of [Rag16] admits a competitive ratio of $O(\log n)$, which is optimal even for randomized algorithms, and [GK19] show that the competitive ratio of the Greedy algorithm lies between $\Omega(n^{0.22})$ and $O(n)$.
For the line metric with adversarial servers and requests, [KP93] show that the competitive ratio of Greedy is at least $12^{n}-$ 1, and a deterministic algorithm with a sublinear competitive ratio is known [ABN+19]. Furthermore, [NR17] present an $O(\log^{2}n)$-competitive deterministic algorithm, which is later shown to be $O(\log n)$-competitive [Rag18]. Moving to randomized algorithms, [GL12] propose an $O(\log n)$-competitive algorithm for the more general doubling metrics. For lower bounds, [FHK05] show that no deterministic algorithm can achieve a competitive ratio strictly better than $9.001$; it is then improved to $\Omega(\sqrt{\log n})$, which holds even for randomized algorithms [PS23].
Stochastic Online Metric Matching.
The stochastic online metric matching problem, where the requests are independently drawn from a known probability distribution, has attracted considerable attention in the past few years. We start with the references for balanced markets. For the Euclidean space $[0,1]^{d}$, where $d$ is a constant not scaling with $n$, [Kan22] proposes the Hierarchical Greedy algorithm with a constant competitive ratio when all servers and requests are uniformly distributed, and [CKKZ23] achieve nearly optimal regret guarantees when servers and requests are allowed to follow different distributions. [BFP23] show that Greedy is $O(1)$-competitive when servers and requests are uniformly distributed in $[0,1]$.
Motivated by the potential value of excess servers, [AALS22] initiate the study of unbalanced markets. When servers and requests are uniformly distributed in $[0,1]$, they show that the total matching cost of Greedy can be substantially improved by transitioning from balanced markets to unbalanced markets, and their results imply that Greedy is $O(\log^{3}n)$-competitive in unbalanced markets. [BFP23] subsequently pinpoint the competitive ratio of Greedy to be $O(1)$ under the same setting. Moreover, Hierarchical Greedy also admits a constant competitive ratio in unbalanced markets when servers and requests are uniformly distributed in $[0,1]^{d}$ for all $d\geq 1$ [Kan22].
In contrast to the fruitful achievements when both servers and requests are stochastic, much less results are known for adversarial servers and stochastic requests. In balanced markets, [GGPW19] give an $O((\log\log\log n)^{2})$-competitive algorithm for general metrics and distributions, which further achieves a constant competitive ratio for tree metrics. Moreover, in balanced markets with requests uniformly distributed in $[0,1]$, [BFP23] show that Greedy is $\Theta(\log n)$-competitive and Hierarchical Greedy is $\Omega(n^{1/4})$-competitive.
Various papers study other aspects of stochastic online metric matching. [TTC94] show that Greedy admits a competitive ratio of $O(\sqrt{n})$ on the unit disk. In addition, [SNL+23] characterize the cost of Greedy in unbalanced markets with servers and requests uniformly distributed in $[0,1]$ using a partial differential equation approach.
Online Algorithm with Predictions.
In recent years, due to the unparalleled impact of machine learning, the paradigm of designing online algorithms using externally provided predictions is of particular interest to overcome both information-theoretic and computational barriers. This paradigm suitably models the practical scenarios where future uncertainty is predictable according to prior knowledge, where predictions can be generated by machine learning models, human experts, historical data, heuristics, or even crowdsourcing. In particular, this line of literature usually assumes the provided predictions are untrusted with their error oblivious to algorithms, and algorithms are required to be consistent and robust. Besides, the performance of algorithms usually deteriorates smoothly as a function of the prediction error when the prediction accuracy decreases.
This paradigm has been incorporated into the study of various classical problems including online metric matching [ACE+23b], metrical task systems [ACE+23b, ACE+23a], online facility location [JLL+22, ACL+21], online bipartite matching [AGKK23, JM22], and caching [GPSS22, AGKP22, LV21], among others.
It is worth of mentioning that [ACE+23b] also study online metric matching with predictions. As the primary difference, we assume that the predictions are request locations and given at the beginning, whereas in their model, each arriving request is accompanied by a prediction of the behaviors of the optimal offline algorithm until the end of the current round, i.e., all servers matched by the optimal offline algorithm for the requests arrived so far. [^2] They allow the predictions to be inconsistent, i.e., each prediction may not be a superset of previous predictions. Moreover, their prediction error is defined against an optimal offline algorithm as the summation of prediction errors of all rounds, where the prediction error of each round is defined as the minimum cost of moving the predicted matched servers to the actual servers matched by the optimal offline algorithm until the end of this round. Hence, their results are in general incomparable with ours.
We note that concurrent and independent work by [CCM25] gives a consistent (but non-robust) algorithm for online metric matching with predictions that is very similar to ours. In particular, they assume that the prediction is given as a distribution over the metric space, and their algorithm draws $n$ independent samples from this distribution as predicted request locations, whereas we directly assume that $n$ points in the metric space are given as a prediction.
2 Model Setup
In our model, $n$ requests and $m$ servers are located in a metric space $(X,\delta)$, where $n\leq m$ and $\delta:X\times X\to{\mathbb{R}}_{\geq 0}$ is a cost function satisfying the triangle inequality. Denote the locations of servers as $S=\{s_{1},\ldots,s_{m}\}$ and the locations of requests as $R=\{r_{1},\ldots,r_{n}\}$. The servers are all available at the beginning with their locations known, and the requests arrive sequentially with their locations revealed only upon arrivals. When a new request arrives, we need to match it to an available server immediately and irrevocably, and then the matched server becomes unavailable. The cost of matching a server $s$ and a request $r$ is $\delta(s,r)$, and we aim to minimize the total matching cost.
Depending on the context, we will assume the locations of servers and requests are either adversarially chosen or independently drawn from a known distribution $\mathbb{D}$ over $X$. For all algorithm ${\mathcal{A}}$, set of servers $S$ with $|S|\geq n$, and set of requests $R$ with $|R|=n$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,R)$ as the cost of ${\mathcal{A}}$ with the initial set of servers being $S$ and the set of requests being $R$; denote $\mathrm{cost}_{{\mathcal{A}}}(S,n)$ as the expected cost of ${\mathcal{A}}$ when the initial set of servers is $S$ and $n$ requests are randomly drawn from $\mathbb{D}$; denote $\mathrm{cost}_{{\mathcal{A}}}(n)$ as the expected cost of ${\mathcal{A}}$ when $|S|=|R|=n$ and all servers and requests are randomly drawn from $\mathbb{D}$.
Given a matching $M$ between two sets $S$ and $R$, for each matched element $s\in S$ (resp. $r\in R$), we use $M(s)$ (resp. $M(r)$) to denote the element in $R$ (resp. $S$) that $s$ (resp. $r$) is matched to. We define the smoothness property of distributions as follows, which requires distributions to admit an upper bounded density function.
Definition 1 ($\beta$-smoothness).
We say that a distribution $\mathbb{D}$ over $X$ is $\beta$-smooth if $\mathbb{D}$ admits a density function $\mu_{\mathbb{D}}$ such that $\mu_{\mathbb{D}}(x)\leq\beta$ for every $x\in X$.
2.1 Benchmark
We adopt the optimal offline cost, defined as the optimal matching cost without any future uncertainty, as our benchmark. For all set of servers $S$ and set of requests $R$ such that $|S|\geq|R|$, denote $\mathrm{OPT}(S,R)$ as the minimum cost of all matchings between $S$ and $R$ that match all requests. Given a set of servers $S$ with $|S|\geq n$ and a distribution $\mathbb{D}$ over $X$, the optimal offline cost on $S$ with $n$ requests randomly drawn from $\mathbb{D}$ is defined as $\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$, where $R\sim\mathbb{D}^{n}$ denotes that the requests in $R$ are independently drawn from $\mathbb{D}$. Furthermore, define $\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]$.
We use competitive ratio and regret to measure the performance of online algorithms. When servers and requests are adversarial, for given metric space $(X,\delta)$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances [^3] We require $|S|=n$ when considering balanced markets and $n\leq|S|\leq O(n)$ for unbalanced markets. $(S,R)$, $\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\alpha\cdot\mathrm{OPT}(S,R)$. Similarly, when servers are adversarial and requests are stochastic, for given metric space $(X,\delta)$ and distribution $\mathbb{D}$, we say that the competitive ratio of an algorithm ${\mathcal{A}}$ is $\alpha$ if for all concerned instances $(S,n)$, $\mathrm{cost}_{{\mathcal{A}}}(S,n)\leq\alpha\cdot\mathrm{OPT}(S,n)$; moreover, define the regret of an algorithm ${\mathcal{A}}$ for a concerned instance $(S,n)$ as
$\mathrm{Regret}({\mathcal{A}};S,n)\triangleq\frac{1}{n}\left(\mathrm{cost}_{{\mathcal{A}}}(S,n)-\mathrm{OPT}(S,n)\right)\,.$
2.2 Optimal Offline Costs for Euclidean Space
A considerable portion of this paper focuses on Euclidean space. In this subsection, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and bound the optimal offline costs for smooth distributions and general distributions. All the omitted proofs in this subsection can be found in Appendix A.
For an arbitrary distribution $\mathbb{D}$, the following upper bounds for $\mathrm{OPT}(n)$ are known from the optimal transport literature.
Lemma 1 (Theorem 3.11 in [HSM22]).
For all $d\geq 1$ and distribution $\mathbb{D}$ on $[0,1]^{d}$,
$\mathrm{OPT}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n}\log n)&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
When restricted to smooth distributions, we improve the upper bound for $d=2$ in Lemma 1 by generalizing the proof of the upper bound in the Ajtai-Komlós-Tusnády Matching Theorem [AKT84], which asserts that $\mathrm{OPT}(n)=\Theta(\sqrt{n\log n})$ for uniform distribution $\mathbb{D}$ and $d=2$.
Lemma 2.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$,
$\mathrm{OPT}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Next, we establish lower bounds for smooth distributions in balanced markets.
Lemma 3.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$,
$\mathrm{OPT}(n)\geq\begin{cases}\Omega(\sqrt{n})&d=1\\ \Omega(n^{1-\frac{1}{d}})&d\geq 2\end{cases}\,.$
The lower bound in Lemma 3 for $d=1$ follows from the quantile mapping and the well-known $\Omega(\sqrt{n})$ lower bound for the optimal offline cost with uniform distribution, and the lower bounds for $d\geq 2$ rely on nearest-neighbor-distance.
Finally, we present lower bounds for smooth distributions in unbalanced markets, which are proved by employing nearest-neighbor-distance as well.
Lemma 4.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $n\leq|S|\leq O(n)$, $\mathrm{OPT}(S,n)\geq\Omega(n^{1-\frac{1}{d}})$.
2.3 The Simulate-Optimize-Assign-Repeat Algorithm
The Simulate-Optimize-Assign-Repeat (SOAR) algorithm is proposed by [CKKZ23] under the setting in which both servers and requests are stochastic. The SOAR algorithm, whose guarantee is stated below, will serve as an important building block of our algorithms for stochastic online metric matching and be invoked in a black-box manner.
Lemma 5 ([CKKZ23]).
Suppose that $m=n$ and $\mathbb{D}$ is a distribution on $X$, where the cost function $\delta$ is not required to satisfy the triangle inequality. When all servers and requests are independently drawn from $\mathbb{D}$, an algorithm exists with an expected cost of $\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)$.
3 A Generic Approach for Beyond the Worst-Case Analysis
In this section, we present a generic algorithmic approach for online metric matching to go beyond the worst-case analysis. Then, we will demonstrate how to apply this approach to leverage the provided distributional information or predicted request locations.
Theorem 1.
Fix a metric space $(X,\delta)$ and an algorithm ${\mathcal{A}}$ for online metric matching. Then, there exists an algorithm ${\mathcal{A}}^{\prime}$ such that for all set of points $I$ given before the arrival of any request, set of servers $S$, and set of requests $R$ with $|S|\geq|I|=|R|$, $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)$. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
Proof.
We construct the algorithm ${\mathcal{A}}^{\prime}$ as follows. Fix three sets of points $S,I,R$ as required. Let $M^{S,I}$ be the matching between $S$ and $I$ with the minimum cost that matches all points in $I$. Then, we run ${\mathcal{A}}$ with the set of initial servers being $I$. For each new request $r$, suppose that ${\mathcal{A}}$ matches it to $i\in I$, and ${\mathcal{A}}^{\prime}$ will match $r$ to $M^{S,I}(i)\in S$.
Since ${\mathcal{A}}^{\prime}$ does not use any randomness except for the invocation of ${\mathcal{A}}$, ${\mathcal{A}}^{\prime}$ is deterministic provided that ${\mathcal{A}}$ is deterministic. By the triangle inequality, the cost incurred by ${\mathcal{A}}^{\prime}$ is upper bounded by
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\sum_{i\in I}\delta(M^{S,I}(i),i)+\mathrm{cost}_{{\mathcal{A}}}(I,R)=\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\,,$
concluding the proof. ∎
3.1 Stochastic Setting
We first apply Theorem 1 to the case where all request locations are independently drawn from a known distribution $\mathbb{D}$. We establish in the following theorem that an algorithm for stochastic servers and requests can be transformed into an algorithm for adversarial servers and stochastic requests in a black-box manner. In particular, for balanced markets, this transformation preserves the competitive ratio of the algorithm up to a constant factor.
Theorem 2.
Given a metric space $(X,\delta)$ and a distribution $\mathbb{D}$ over $X$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)$.
The rest of this subsection is devoted to proving Theorem 2. We start with proving the triangle inequality for $\mathrm{OPT}(\cdot,\cdot)$.
Lemma 6.
For all sets of points $S,I,R$ with $|S|=|I|=|R|$, $\mathrm{OPT}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{OPT}(I,R)$.
Proof.
Given $S,I,R$ with $|S|=|I|=|R|$, let $M_{1}$ be the min-cost perfect matching between $S$ and $I$, and let $M_{2}$ be the min-cost perfect matching between $I$ and $R$. Note that
$\mathrm{OPT}(S,I)+\mathrm{OPT}(I,R)=\sum_{i\in I}\delta(M_{1}(i),i)+\sum_{i\in I}\delta(i,M_{2}(i))\geq\sum_{i\in I}\delta(M_{1}(i),M_{2}(i))\geq\mathrm{OPT}(S,R)\,,$
where the first inequality holds by the triangle inequality. ∎
Assuming that requests are stochastic, we start by lower bounding the optimal offline cost in balanced markets with adversarial servers. In particular, we show that it is at least the optimal offline cost in balanced markets with stochastic servers and requests up to a constant factor.
Lemma 7.
For all metric space $(X,\delta)$, distribution $\mathbb{D}$ over $X$, and set of servers $S$ with $|S|=n$, $\mathrm{OPT}(S,n)\geq\frac{1}{2}\cdot\mathrm{OPT}(n)$.
Proof.
Given a set of servers $S$ with $|S|=n$, by Lemma 6, we know that $\mathrm{OPT}(I,R)\leq\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)$ for all sets of points $I,R$ with $|I|=|R|=n$. Taking expectations on both sides over $I,R\sim\mathbb{D}^{n}$, we obtain
$\mathrm{OPT}(n)\leq\mathbb{E}_{I,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(I,S)+\mathrm{OPT}(S,R)\right]=2\cdot\mathrm{OPT}(S,n)\,,$
which concludes the proof. ∎
Now, we are ready to prove Theorem 2. Given a metric space $(X,\delta)$ and a distribution $\mathbb{D}$ over $X$, let ${\mathcal{A}}$ be an algorithm such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$, and we construct an algorithm ${\mathcal{A}}^{\prime}$ for adversarial servers and stochastic requests as follows. Given as input a set of servers $S$ and a set of requests $R$ with $|S|\geq|R|=n$, we randomly generate a set of points $I\sim\mathbb{D}^{n}$ before the arrival of any request, and let ${\mathcal{A}}^{\prime}$ be the algorithm as per Theorem 2, which follows that
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\leq\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\,.$
Since $I$ and $R$ are independent, the expected cost of ${\mathcal{A}}^{\prime}$ is upper bounded by
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathbb{E}_{I,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,I)+\mathrm{cost}_{{\mathcal{A}}}(I,R)\right]=\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)\,,$
which concludes the first part in Theorem 2. Furthermore, when $|S|=n$, by the guarantee of ${\mathcal{A}}$ and Lemma 7, we obtain
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\alpha\cdot\mathrm{OPT}(n)\leq(2\alpha+1)\cdot\mathrm{OPT}(S,n)\,,$
concluding the second part in Theorem 2.
3.2 Prediction Setting
In this subsection, we focus on balanced markets, i.e., $m=n$, and assume that the locations of servers and requests are adversarially chosen. In addition, we assume that algorithms are equipped with a prediction of request locations. We then apply Theorem 1 to give an algorithm whose performance smoothly depends on the prediction error, i.e., a consistent algorithm. More formally, the prediction is given in the form of a set of predicted request locations $R^{\prime}$ with $|R^{\prime}|=n$, which is provided before the arrival of any request, and the prediction error is defined as $\eta\triangleq\mathrm{OPT}(R^{\prime},R)$. We reiterate that the prediction error is not known by algorithms. Throughout, we slightly abuse notations and use $\mathrm{OPT}\triangleq\mathrm{OPT}(S,R)$ to denote the optimal offline cost.
Theorem 3.
Given a metric space $(X,\delta)$, assume that there exists an $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions. Then, there exists an algorithm ${\mathcal{A}}^{\prime}$ such that, when equipped with a prediction of request locations with error $\eta$, ${\mathcal{A}}^{\prime}$ is $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
Proof.
We construct ${\mathcal{A}}^{\prime}$ as follows. Given as input a set of servers $S$, a set of requests $R$, and a prediction of request locations $R^{\prime}$ with $|S|=|R|=|R^{\prime}|$ and $\eta=\mathrm{OPT}(R,R^{\prime})$, let ${\mathcal{A}}^{\prime}$ be the algorithm as per Theorem 2 by setting $I=R^{\prime}$. It gives that,
$\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)$ $\leq\mathrm{OPT}(S,R^{\prime})+\mathrm{cost}_{{\mathcal{A}}}(R^{\prime},R)\leq\mathrm{OPT}(S,R)+\mathrm{OPT}(R,R^{\prime})+\mathrm{cost}_{{\mathcal{A}}}(R^{\prime},R)$
$\leq\mathrm{OPT}(S,R)+\mathrm{OPT}(R,R^{\prime})+\alpha\cdot\mathrm{OPT}(R^{\prime},R)=\mathrm{OPT}(S,R)+\eta(\alpha+1)\,,$
where the first inequality holds by the guarantee of Theorem 2, the second inequality holds by Lemma 6, and the third inequality holds due to the competitive ratio of ${\mathcal{A}}$. Finally, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ being deterministic follows from the guarantee of Theorem 2. ∎
4 Stochastic Setting: Euclidean Space with Smooth Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance and $\mathbb{D}$ to be $\beta$-smooth [^4] The smoothness of distributions is a mild technical assumption allowing us to lower bound $\mathrm{OPT}(n)$. To see the necessity, consider $\mathbb{D}$ a non-smooth distribution supported on a singleton, in which case $\mathrm{OPT}(n)=0$. for some constant $\beta>0$, and we will apply Theorem 2 to derive algorithms for both balanced and unbalanced markets. In Section 4.1, we focus on uniform distribution, which is a special case of smooth distributions, and give an $O(1)$-competitive algorithm for all $d\geq 1$ in balanced markets.
Next, we consider balanced markets with smooth distributions in Section 4.2. Recall that when $d=1$, an $O(1)$-competitive algorithm in balanced markets for general distributions is given by [GGPW19]. Moreover, this algorithm is $O((\log\log\log n)^{2})$-competitive for general metrics and distributions. However, for $d=2$ and smooth distributions, we are only able to estimate the optimal offline cost up to a $\Theta(\sqrt{\log n})$ factor, which prevents our approach from achieving an $o(\sqrt{\log n})$ competitive ratio. Hence, we only focus on the case with $d\geq 3$, where we present an $O(1)$-competitive algorithm.
Finally, in Section 4.3, we extend our results for smooth distributions to unbalanced markets that possess at most linearly many excess servers, i.e., $n\leq m\leq O(n)$. We give an algorithm with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
4.1 Prelude: Balanced Markets with Uniform Distribution
In this subsection, we consider balanced markets and assume $\mathbb{D}$ to be uniform distribution. In the following theorem, we show that combining Theorem 2 with the Hierarchical Greedy algorithm of [Kan22] results in an $O(1)$-competitive algorithm for all $d\geq 1$.
Theorem 4.
Suppose that $m=n$. For $d\geq 1$, an $O(1)$-competitive algorithm exists for adversarial servers and uniformly distributed requests.
Proof.
Recall that [Kan22] gives an $O(1)$-competitive algorithm for $d\geq 1$ when both servers and requests follow uniform distribution. Combining the above algorithm with Theorem 2 concludes the proof. ∎
4.2 Balanced Markets
In this subsection, we assume $d\geq 3$ and consider balanced markets with smooth distributions. We will prove that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ when servers and requests are randomly drawn from $\mathbb{D}$. Together with Theorem 2, this leads to an $O(1)$-competitive algorithm for $d\geq 3$ with adversarial servers and stochastic requests.
We first present our main theorem in this subsection.
Theorem 5.
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$, an $O(1)$-competitive algorithm exists for $d\geq 3$.
The rest of this subsection is devoted to proving Theorem 5. By Theorem 2, it suffices to show that the SOAR algorithm is $O(1)$-competitive for $d\geq 3$ when both servers and requests are randomly drawn from $\mathbb{D}$. We compute the expected cost of the SOAR algorithm in the following lemma. Even though we only consider $d\geq 3$ in this subsection, the costs of the SOAR algorithm for $d\leq 2$ computed by Lemma 8 will be useful afterward.
Lemma 8.
Suppose that $m=n$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. An algorithm ${\mathcal{A}}$ exists such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n\log n})&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Proof.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5, and we estimate the cost of ${\mathcal{A}}$ for $d=1$, $d=2$, and $d\geq 3$ separately. Recall that an upper bound for $\mathrm{OPT}(n)$ is given by Lemma 2. For $d=1$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n}\right)\,.$
For $d=2$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\sqrt{\frac{\log t}{t}}\right)=O\left(\sqrt{\log n}\cdot\int_{1}^{n}\frac{1}{\sqrt{x}}\mathrm{d}x\right)=O\left(\sqrt{n\log n}\right)\,.$
For $d\geq 3$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{t^{1/d}}\right)=O\left(\int_{1}^{n}\frac{1}{x^{1/d}}\mathrm{d}x\right)=O\left(n^{1-\frac{1}{d}}\right)\,.$
Combining all the above concludes the proof. ∎
Finally, combining Lemma 8 with the lower bounds for $\mathrm{OPT}(n)$ given by Lemma 3 concludes the proof of Theorem 5.
4.3 Unbalanced Markets
We extend our results for smooth distributions to unbalanced markets satisfying $n\leq m\leq O(n)$.
Theorem 6.
Suppose that $n\leq m\leq O(n)$ and $\mathbb{D}$ is $\beta$-smooth for some constant $\beta>0$. When servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$, an algorithm exists with a competitive ratio of $O(\sqrt{n})$ for $d=1$, $O(\sqrt{\log n})$ for $d=2$, and $O(1)$ for $d\geq 3$.
Proof.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 8. By Theorem 2, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $n\leq|S|\leq O(n)$. Combining the lower bounds for $\mathrm{OPT}(S,n)$ given by Lemma 4 concludes the proof. ∎
5 Stochastic Setting: Euclidean Space with General Distributions
In this section, we assume the metric space to be $[0,1]^{d}$ with the Euclidean distance. We do not make any assumptions on the distribution and achieve nearly optimal regret bounds for all $d\geq 1$ in unbalanced markets with an arbitrary number of excess servers. In particular, our results recover the nearly optimal regret bounds by [CKKZ23], which is established in balanced markets with servers and requests following different distributions. In the weaker setting studied by [CKKZ23], the regret bound is optimal for $d\neq 2$ and optimal up to a logarithmic factor for $d=2$, implying that our regret bounds are also nearly optimal.
Theorem 7.
Suppose that $n\leq m$ and $\mathbb{D}$ is an arbitrary distribution over $[0,1]^{d}$. An algorithm ${\mathcal{A}}^{\prime}$ exists that, for every set of servers $S$ with $|S|=m$, admits a regret of
$\mathrm{Regret}({\mathcal{A}}^{\prime};S,n)\leq\begin{cases}O(n^{-\frac{1}{2}})&d=1\\ O(n^{-\frac{1}{2}}\log n)&d=2\\ O(n^{-\frac{1}{d}})&d\geq 3\end{cases}\,.$
Proof.
For an arbitrary distribution $\mathbb{D}$, by applying the same arguments as in the proof of Lemma 8 except that the upper bound for $\mathrm{OPT}(n)$ is given by Lemma 1 instead of Lemma 2, we obtain an algorithm ${\mathcal{A}}$ such that
$\mathrm{cost}_{{\mathcal{A}}}(n)\leq\begin{cases}O(\sqrt{n})&d=1\\ O(\sqrt{n}\log n)&d=2\\ O(n^{1-\frac{1}{d}})&d\geq 3\end{cases}\,.$
By Theorem 2, an algorithm ${\mathcal{A}}^{\prime}$ exists satisfying $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n)$ for every set of servers $S$ with $|S|\geq n$. As a result, the regret of ${\mathcal{A}}^{\prime}$ equals
$\mathrm{Regret}({\mathcal{A}}^{\prime};S,n)=\frac{1}{n}\left(\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)-\mathrm{OPT}(S,n)\right)\leq\frac{\mathrm{cost}_{{\mathcal{A}}}(n)}{n}\,,$
which concludes the proof. ∎
We briefly explain the reason behind the contrast that in unbalanced markets with $d\leq 2$, our approach admits a nearly optimal regret (Theorem 7) with a suboptimal competitive ratio (Theorem 6). In fact, our regret guarantee indicates that $\mathrm{cost}_{{\mathcal{A}}}(S,n)-\mathrm{OPT}(S,n)$ is small, yet we have no control over the lower bound for $\mathrm{OPT}(S,n)$ even with smooth distributions, which prevents us from achieving a nearly optimal competitive ratio.
6 Prediction Setting
In this section, we adopt the same setting and notations as in Section 3.2. We will present an algorithm with desirable consistency and robustness guarantees, whose performance also smoothly depends on the prediction error. Recall that $\eta$ denotes the prediction error and $\mathrm{OPT}$ denotes the optimal offline cost, and our main result in this section is presented as follows.
Theorem 8.
Assume that $m=n$. Given an algorithm ${\mathcal{A}}$ that does not use predictions and is $\alpha$-competitive for a particular metric, there exists an algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions and is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive for the same metric. Moreover, if ${\mathcal{A}}$ is deterministic, then ${\mathcal{A}}^{\prime}$ is also deterministic.
The proof of Theorem 8 is deferred to Section 6.1, and we now apply Theorem 8 to the existing algorithms in the literature. Recall that without predictions, there exist deterministic $(2n-1)$-competitive algorithms for general metrics [KP93, KMV94, Rag16], a deterministic $O(\log n)$-competitive algorithm for the line metric [Rag18], a randomized $O(\log^{2}n)$-competitive algorithm for general metrics [BBGN14], and a randomized $O(\log n)$-competitive algorithm for doubling metrics [GL12]. As a corollary, we derive the following results.
Corollary 1.
For general metrics, there exist a deterministic $O(\min\{n,1+\frac{\eta\cdot n}{\mathrm{OPT}})\}$-competitive algorithm and a randomized $O(\min\{\log^{2}n,1+\frac{\eta\cdot\log^{2}n}{\mathrm{OPT}}\})$-competitive algorithm. For the line metric, there exists a deterministic $O(\min\{\log n,1+\frac{\eta\cdot\log n}{\mathrm{OPT}})\}$-competitive algorithm. For doubling metrics, there exists a randomized $O(\min\{\log n,1+\frac{\eta\cdot\log n}{\mathrm{OPT}})\}$-competitive algorithm.
6.1 Proof of Theorem 8
Recall that we have given a non-robust algorithm with the desired consistency guarantee in Theorem 3, and now we show how to robustify it, which leads to the desired algorithm.
A powerful machinery for deterministically combining several algorithms into one in a black-box manner is proposed by [FRR94] for the $k$-server problem and is subsequently applied by [LV21, ACE+23b] under the context of online algorithm with predictions. We will use the same approach to combine our consistent but non-robust algorithm with the given algorithm ${\mathcal{A}}$ that does not use predictions, which is by definition robust but inconsistent. Consequently, robustness and consistency are simultaneously achieved.
We adapted the framework for combining several online algorithms generalized by [ACE+23b] to online metric matching, which is a direct consequence of combining [ACE+23b, Theorem 18] and [ACE+23b, Lemma 21].
Lemma 9 (Theorem $18$ and Lemma $21$ in [ACE+23b]).
Given $k$ deterministic algorithms ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$ for online metric matching, for every $1<\gamma\leq 2$, there exists a deterministic algorithm ${\mathcal{A}}^{\mathrm{com}}$ for online metric matching such that for all $S$ and $R$ with $|S|=|R|$,
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,.$
We generalize Lemma 9 to combine randomized online algorithms in the following lemma.
Lemma 10.
Given $k$ randomized algorithms ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$ for online metric matching, for every $1<\gamma\leq 2$, there exists a randomized algorithm ${\mathcal{A}}^{\mathrm{com}}$ for online metric matching such that for all $S$ and $R$ with $|S|=|R|$,
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,.$
Proof.
For each ${\mathcal{A}}_{i}$, recall that ${\mathcal{A}}_{i}$ is randomized, and we use ${\mathcal{A}}_{i}(S,R\mid B)$ to denote the cost incurred by ${\mathcal{A}}_{i}$ with servers $S$ and requests $R$ conditional on the random bits used by ${\mathcal{A}}_{i}$ being $B=b_{1}b_{2}\cdots$. To construct our desired algorithm ${\mathcal{A}}^{\mathrm{com}}$, we first independently sample the random bits $B=b_{1}b_{2}\cdots$ used by ${\mathcal{A}}_{1},\ldots,A_{k}$ and then apply the combination scheme of Lemma 9 to combine ${\mathcal{A}}_{1},\ldots,{\mathcal{A}}_{k}$. By the guarantee of Lemma 9, for all $S$ and $R$, it holds that
$\mathrm{cost}_{{\mathcal{A}}^{\mathrm{com}}}(S,R)$ $\leq\mathbb{E}_{B}\left[\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R\mid B)\}\right]$
$\leq\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathbb{E}_{B}[\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R\mid B)]\}$
$=\left(\frac{2\gamma^{k}}{\gamma-1}+1\right)\cdot\min_{i\in[k]}\{\mathrm{cost}_{{\mathcal{A}}_{i}}(S,R)\}\,,$
where the second inequality follows from Jensen’s inequality. ∎
Now, we are ready to prove Theorem 8. We only prove the statement for deterministic algorithms by using Theorem 3 and Lemma 9, and the statement for randomized algorithms can be established analogously by applying Theorem 3 and Lemma 10. Given a deterministic $\alpha$-competitive algorithm ${\mathcal{A}}$ that does not use predictions for a particular metric, by Theorem 3, there exists a deterministic $(1+\frac{\eta(\alpha+1)}{\mathrm{OPT}})$-competitive algorithm ${\mathcal{A}}^{\prime}$ that leverages predictions for the same metric. By applying Lemma 9 with $\gamma=2$ to combine ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, we obtain a deterministic algorithm ${\mathcal{A}}^{\mathrm{com}}$ incurring a cost of $O(\min\{\mathrm{cost}_{{\mathcal{A}}}(S,R),\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,R)\})$ for all $S$ and $R$ with $|S|=|R|$, which implies that ${\mathcal{A}}^{\mathrm{com}}$ is $O(\min\{\alpha,1+\frac{\eta(\alpha+1)}{\mathrm{OPT}}\})$-competitive by the competitive ratios of ${\mathcal{A}}$ and ${\mathcal{A}}^{\prime}$, concluding the proof.
7 Adversarial Setting
In this section, we focus on a more general setting where both servers and requests are adversarially chosen. This setting has been extensively studied in the past [Rag18, BBGN14], whereas all known guarantees are established in balanced markets. We show that the competitive ratio guarantees in balanced markets of algorithms satisfying certain natural conditions can be carried over to unbalanced markets, and the resulting competitive ratios in unbalanced markets only depend on the number of requests. Combined with the algorithms of prior work, we obtain competitive ratio guarantees in unbalanced markets with adversarial servers and requests, which in turn leads to improved guarantees in unbalanced markets with adversarial servers and stochastic requests. In particular, we show that both the $O(\log^{3}n)$-competitive algorithm of [MNP06] for general metrics and the $O(\log n)$-competitive algorithm of [GL12] for doubling metrics admit the same competitive ratios in unbalanced markets. The latter result also improves the $O(\sqrt{n})$ competitive ratio for $d=1$ in unbalanced markets with servers adversarially chosen and requests drawn from a smooth distribution in Theorem 6.
Given the execution of an online algorithm ${\mathcal{A}}$ on an instance with initial servers $S$, we use $N_{{\mathcal{A}}}(r\mid S)$ to denote the set of available servers closest to the request $r$ upon its arrival. We say that an online algorithm ${\mathcal{A}}$ is greedy if it always (randomly) matches a new request to one of the closest available servers, i.e., it matches $r$ to a server in $N_{{\mathcal{A}}}(r\mid S)$. In particular, every greedy algorithm is completely determined by a (randomized) tie-breaking rule such that when facing multiple closest available servers upon the arrival of a new request, the algorithm applies the tie-breaking rule to (randomly) decide who to match with. We say that the tie-breaking rule of a greedy algorithm ${\mathcal{A}}$ is local if it only depends on the set of closest available servers $N_{{\mathcal{A}}}(r\mid S)$. Moreover, the local tie-breaking rule of a greedy algorithm ${\mathcal{A}}$ is called non-increasing if the probability of selecting each server in $N_{{\mathcal{A}}}(r\mid S)$ does not increase by adding additional servers into $N_{{\mathcal{A}}}(r\mid S)$, i.e., for all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,.$
We show that the competitive ratio guarantee for every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets can be carried over to unbalanced markets.
Theorem 9.
For an arbitrary metric, if a greedy algorithm ${\mathcal{A}}$ with a local and non-increasing tie-breaking rule is $\alpha$-competitive in balanced markets with adversarial servers and requests, then ${\mathcal{A}}$ is $\alpha$-competitive in unbalanced markets with adversarial servers and requests.
We remark here that the resulting competitive ratio $\alpha$ for unbalanced markets can depend on the number of requests but is independent of the number of excess servers.
Proof of Theorem 9.
For two sets of points $S$ and $R$ in the metric with $|S|\geq\mid R|$, denote $\mathrm{cost}_{{\mathcal{A}}}(S,R)$ as the expected cost of ${\mathcal{A}}$ when the set of initial servers is $S$ and the set of requests is $R$. Then, we claim that for all sets of points $S,S^{\prime},R$ in the metric with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, $\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$.
Before proving our claim, we first show the desired competitive ratio guarantee of ${\mathcal{A}}$ in unbalanced markets by applying the claim. Let $S$ be the set of initial servers and $R$ be the set of requests with $|S|\geq|R|=n$, and let $S^{\prime}\subseteq S$ satisfying $|S^{\prime}|=n$ and $\mathrm{OPT}(S,R)=\mathrm{OPT}(S^{\prime},R)$, i.e., $S^{\prime}$ consists of all matched servers in $S$ in the minimum cost matching between $S$ and $R$ that matches all requests in $R$. Since ${\mathcal{A}}$ is $\alpha$-competitive in balanced markets,
$\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S^{\prime},R)=\alpha\cdot\mathrm{OPT}(S,R)\,.$
Given that $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|$, by our claim,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\leq\alpha\cdot\mathrm{OPT}(S,R)\,,$
implying that ${\mathcal{A}}$ is $\alpha$-competitive in unbalanced markets.
It remains to prove our claim, which we establish by induction on $|R|$. When $|R|=0$, we have $\mathrm{cost}_{{\mathcal{A}}}(S,R)=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)=0$, and the statement straightforwardly holds. Assume for induction that the statement holds for $|R|=k$, and we show that the statement also holds for $|R|=k+1$. Given $S,S^{\prime},R$ with $S^{\prime}\subseteq S$ and $|S^{\prime}|\geq|R|=k+1$, let $r\in R$ be the first arrived request. We use $M_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})$ to denote the (random) server that ${\mathcal{A}}$ matches $r$ to when the set of initial servers being $S$ and $S^{\prime}$, respectively. Since ${\mathcal{A}}$ is greedy, $M_{{\mathcal{A}}}(r\mid S)\in N_{{\mathcal{A}}}(r\mid S)$ and $M_{{\mathcal{A}}}(r\mid S^{\prime})\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. The following lemma states that either $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$ or $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$.
Lemma 11.
One of the following must hold: $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$ or $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$.
Proof.
Suppose for contradiction that $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})\neq\emptyset$ and $N_{{\mathcal{A}}}(r\mid S^{\prime})\not\subseteq N_{{\mathcal{A}}}(r\mid S)$, and let $s\in N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\setminus N_{{\mathcal{A}}}(r\mid S)$. Since $r$ is the first arrived request, all servers in $S$ (resp. $S^{\prime}$) are available upon the arrival of $r$, and hence, $N_{{\mathcal{A}}}(r\mid S)$ (resp. $N_{{\mathcal{A}}}(r\mid S^{\prime})$) contains all servers in $S$ (resp. $S^{\prime}$) closest to $r$. Moreover, note that $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq S^{\prime}\subseteq S$. On one hand, since $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, $\delta(s,r)=\delta(s^{\prime},r)$. On the other hand, since $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\notin N_{{\mathcal{A}}}(r\mid S)$, $\delta(s,r)<\delta(s^{\prime},r)$, leading to a contradiction. ∎
We analyze the expected cost of ${\mathcal{A}}$ under both cases in Lemma 11 separately. We start with the easier case where $N_{{\mathcal{A}}}(r\mid S)\cap N_{{\mathcal{A}}}(r\mid S^{\prime})=\emptyset$, which implies that $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$ and $\delta(r,s)<\delta(r,s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. By $N_{{\mathcal{A}}}(r\mid S)\cap S^{\prime}=\emptyset$, we know that $M_{{\mathcal{A}}}(r\mid S)\notin S^{\prime}$ and hence $S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\}\subseteq S^{\prime}\subseteq S\setminus\{M_{{\mathcal{A}}}(r\mid S)\}$. By the inductive hypothesis, $\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})\leq\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})$. As a result,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)$ $=\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S))+\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{M_{{\mathcal{A}}}(r\mid S)\},R\setminus\{r\})]$
$<\mathbb{E}[\delta(r,M_{{\mathcal{A}}}(r\mid S^{\prime}))+\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{M_{{\mathcal{A}}}(r\mid S^{\prime})\},R\setminus\{r\})]$
$=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)\,,$
concluding the proof.
Next, we consider the case where $N_{{\mathcal{A}}}(r\mid S^{\prime})\subseteq N_{{\mathcal{A}}}(r\mid S)$, which implies that $\delta(r,s)=\delta(r,s^{\prime})$ for all $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S)$. For every $s\in N_{{\mathcal{A}}}(r\mid S)$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, define
$P_{S}(s)$ =$=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S)=s\right\},$ $Q_{S}(s)$ $=\mathrm{cost}_{{\mathcal{A}}}(S\setminus\{s\},R\setminus\{r\})\,,$
$P_{S^{\prime}}(s^{\prime})$ =$=\mathbb{P}\left\{M_{{\mathcal{A}}}(r\mid S^{\prime})=s^{\prime}\right\},$ $Q_{S^{\prime}}(s^{\prime})$ $=\mathrm{cost}_{{\mathcal{A}}}(S^{\prime}\setminus\{s^{\prime}\},R\setminus\{r\})\,.$
Since the tie-breaking rule of ${\mathcal{A}}$ is local and non-increasing, $P_{S^{\prime}}(s)\geq P_{S}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. As a result,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)-\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$
$=\sum_{s\in N_{{\mathcal{A}}}(r\mid S)}P_{S}(s)\cdot(\delta(r,s)+Q_{S}(s))-\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S^{\prime}}(s)\cdot(\delta(r,s)+Q_{S^{\prime}}(s))$
$=\sum_{s\in N_{{\mathcal{A}}}(r\mid S)}P_{S}(s)\cdot Q_{S}(s)-\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S^{\prime}}(s)\cdot Q_{S^{\prime}}(s)$
$=\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}(P_{S}(s)\cdot Q_{S}(s)-P_{S^{\prime}}(s)\cdot Q_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)$
$\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s)\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)\,,$
where the second equality holds since $\delta(r,s)=\delta(r,s^{\prime})$ for all $s,s^{\prime}\in N_{{\mathcal{A}}}(r\mid S)$ and the inequality holds since $Q_{S}(s)\leq Q_{S^{\prime}}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$ by the inductive hypothesis. Note that for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$, we have $S^{\prime}\setminus\{s^{\prime}\}\subseteq S^{\prime}\subseteq S\setminus\{s\}$ and hence $Q_{S}(s)\leq Q_{S^{\prime}}(s^{\prime})$ by the inductive hypothesis. Let $B\triangleq\min_{s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s^{\prime})$, and it follows that $Q_{S}(s)\leq B\leq Q_{S^{\prime}}(s^{\prime})$ for all $s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})$ and $s^{\prime}\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. Therefore,
$\mathrm{cost}_{{\mathcal{A}}}(S,R)-\mathrm{cost}_{{\mathcal{A}}}(S^{\prime},R)$ $\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}Q_{S^{\prime}}(s)\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot Q_{S}(s)$
$\leq\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}B\cdot(P_{S}(s)-P_{S^{\prime}}(s))+\sum_{s\in N_{{\mathcal{A}}}(r\mid S)\setminus N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S}(s)\cdot B$
$=B\cdot\left(\sum_{s\in N_{{\mathcal{A}}}(r\mid S)}P_{S}(s)-\sum_{s\in N_{{\mathcal{A}}}(r\mid S^{\prime})}P_{S^{\prime}}(s)\right)$
$=0\,,$
where the second inequality holds because $P_{S^{\prime}}(s)\geq P_{S}(s)$ for every $s\in N_{{\mathcal{A}}}(r\mid S^{\prime})$. This concludes the proof. ∎
7.1 Applications on Hierarchically Separated Tree Metrics
Now, we apply Theorem 9 to algorithms of prior work. Recall that for balanced markets with adversarial servers and requests, the algorithm of [MNP06], abbreviated as the MNP algorithm, is $O(\log^{3}n)$-competitive for general metrics, and the Random-Subtree algorithm of [GL12] is $O(\log n)$-competitive for doubling metrics. We show that the competitive ratios of both algorithms are retained in unbalanced markets.
Both the MNP and Random Subtree algorithms are designed based on the same framework. They first randomly embed the original metric into a Hierarchically Separated Tree (HST) metric, which preserves the distance between every pair of nodes up to an $O(\log n)$ distortion. Next, they show that a randomized greedy subroutine with a specific tie-breaking rule admits a competitive ratio of $\beta$ for the HST metric. Overall, this leads to an $O(\beta\log n)$-competitive algorithm for the original metric. Hence, for our purpose, it suffices to show that the tie-breaking rule of the greedy subroutine for the HST metric is local and non-increasing.
Recall that an HST is a rooted tree satisfying the following property: For each subtree, the distances from the subtree’s root to all leaf nodes in the subtree are identical. Moreover, the metric embedding phase of the algorithms will map each location in the original metric to a leaf node of the HST. From now on, we will restrict our attention to the HST metric with all servers and requests locating at the HST’s leaf nodes.
We start by analyzing the greedy subroutine adopted by the MNP algorithm, which works as follows. Upon the arrival of a new request $r$, assuming the set of initial servers to be $S$, denote $A(r\mid S)$ as the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, and the subroutine matches $r$ to a uniformly random server among all available servers in the subtree rooted at $A(r\mid S)$. In the following lemma, we show that this tie-breaking rule is local and non-increasing, which implies that the greedy subroutine in the MNP algorithm satisfies the conditions in Theorem 9.
Lemma 12.
The greedy subroutine in the MNP algorithm admits a local and non-increasing tie-breaking rule.
The proof of Lemma 12 is deferred to Section B.1. The following corollary is a direct consequence of combining Theorem 9 and Lemma 12 with the competitive ratio guarantee of the MNP algorithm in balanced markets [MNP06].
Corollary 2.
Suppose that $m$ servers and $n$ requests are adversarially chosen with $n\leq m$. For an arbitrary metric, an $O(\log^{3}n)$-competitive algorithm exists.
We remark here that for general metrics with adversarial servers and requests, a randomized algorithm with an improved competitive ratio of $O(\log^{2}n)$ exists [BBGN14]. However, this algorithm does not belong to the family of greedy algorithms, and hence, Theorem 9 does not apply.
Next, we analyze the greedy subroutine adopted by the Random-Subtree algorithm, which works as follows. Upon the arrival of a new request $r$, recall that $A(r\mid S)$ denotes the lowest ancestor of $r$ the subtree rooted at whom contains at least one available server, assuming the set of initial servers to be $S$. Starting off at $A(r\mid S)$, the subroutine repeatedly moves us to a uniformly random subtree among all subtrees of the current node containing at least one available server until we reach a leaf node, and $r$ is matched to a uniformly random available server at the leaf node. In the following lemma, we show that this tie-breaking rule is local and non-increasing, which implies that the greedy subroutine in the Random Subtree algorithm satisfies the conditions in Theorem 9.
Lemma 13.
The greedy subroutine in the Random Subtree algorithm admits a local and non-increasing tie-breaking rule.
The proof of Lemma 13 is deferred to Section B.2. Combining Theorem 9 and Lemma 13 with the competitive ratio guarantee of the Random Subtree algorithm in balanced markets [GL12], we derive the competitive ratio guarantee of the Random Subtree algorithm in unbalanced markets in the following corollary.
Corollary 3.
Suppose that $m$ servers and $n$ requests are adversarially chosen with $n\leq m$. For doubling metrics, an $O(\log n)$-competitive algorithm exists.
8 Discussion and Future Directions
In this paper, we present a generic approach for online metric matching to utilize the provided distributional knowledge or predictions of request locations, leading to new results in both settings.
When requests are stochastic, for balanced markets, we achieve an $O(1)$ competitive ratio for the Euclidean space with smooth distributions and $d\neq 2$. Nevertheless, it still remains open to give an $O(1)$-competitive algorithm for general metrics and distributions. The potential approaches for achieving this goal comprise (1) improving the analysis of the algorithm of [GGPW19], which is proved to be $O((\log\log\log n)^{2})$-competitive without any super-constant lower bound known, and (2) applying Theorem 2 and the SOAR algorithm to the simplified model given by [GGPW19]. The main technical barrier underlying the latter approach lies in the difficulty of characterizing $\mathrm{OPT}(n)$ for general metrics and distributions. In Appendix D, we provide further evidence that supports the plausibility of this approach by applying it to achieve a constant competitive ratio with tree metrics and general distributions, recovering a result achieved by the algorithm of [GGPW19].
We also derive an algorithm in unbalanced markets with improved competitive ratios when requests are drawn from smooth distributions over the Euclidean space. Our algorithm admits an optimal constant competitive ratio for $d\geq 3$, leaving the improvement for $d\leq 2$ an intriguing future direction. Given that the line metric, i.e., $d=1$, is typically the most tractable case [GGPW19, BFP23], it is surprising that it becomes the most intractable one with the presence of excess servers, rendering its resolution more appealing.
Moreover, for adversarial servers and requests without predictions, we show that the competitive ratio of every greedy algorithm with a local and non-increasing tie-breaking rule in balanced markets is preserved in unbalanced markets. However, our characterization fails to capture all the existing algorithms. In this most extensively studied setting, our understanding of the achievable guarantees in unbalanced markets remains very limited, and nothing beyond our results is known. An exciting future direction is to generalize the performance of the existing algorithms with provable guarantees in balanced markets to unbalanced markets.
Several open questions arise from the modeling assumptions in this paper. First, our results assume advance knowledge of $n$, raising the question of whether competitive guarantees are achievable without this information. Next, while we define unbalanced markets as $n\leq m\leq O(n)$ for our competitive ratio results in the stochastic setting, a definition adopted by all prior work, it remains unclear whether these results extend to the cases where $m$ is polynomially or exponentially larger than $n$ [^5] Notice that our results for regret analysis and those in the adversarial setting hold regardless of the number of excess servers. In particular, as $m$ increases, the competitive ratio of our algorithm also increases correspondingly with a context-dependent increasing rate, and when $m=\omega(n)$, our algorithm no longer achieves a constant competitive ratio for the Euclidean space with $d\geq 3$. Moreover, relaxing our assumptions that servers outnumber requests and all requests must be served leads to an interesting variant where request rejection is permitted—a setting that bears similarities to the multi-unit prophet inequality problem [JMZ22]. Finally, it is appealing to relax the assumption that the cost function satisfies the triangle inequality, and we establish positive results for cost functions only approximately satisfying the triangle inequality in Appendix C.
Appendix A Postponed Proofs in Section 2.2
A.1 Proof of Lemma 2
The upper bounds for $d\neq 2$ directly follow from Lemma 1, and it remains to prove the upper bound for $d=2$. We use $\mathbb{U}$ to denote uniform distribution. Recall that the Ajtai-Komlós-Tusnády Matching Theorem [AKT84] states that $\mathbb{E}_{S,R\sim{\mathbb{U}}^{n}}[\mathrm{OPT}(S,R)]=\Theta(\sqrt{n\log n})$ for $d=2$. We will generalize the proof of the upper bound in the Ajtai-Komlós-Tusnády Matching Theorem interpreted by [Tal92, Section 4.3].
We first set up some notations necessary for the proof. For all metric space $(X,\delta)$ and $X^{\prime}\subseteq X$, let $\Delta(X^{\prime})$ denote the diameter of $X^{\prime}$ with respect to $\delta$, i.e.,
$\Delta(X^{\prime})=\sup_{x,y\in X^{\prime}}\delta\left(x,y\right)\,.$ (1)
Given a set $X$, an admissible sequence of $X$ is an increasing sequence $({\mathcal{A}}_{t})_{t\geq 0}$ of partitions of $X$ where the cardinality of ${\mathcal{A}}_{t}$ satisfies $|{\mathcal{A}}_{0}|=1$ and $|{\mathcal{A}}_{t}|\leq 2^{2^{t}}$ for $t\geq 1$. Given an admissible sequence $({\mathcal{A}}_{t})_{t\geq 0}$ of $X$, we use $A_{t}(x)$ to denote the unique element of ${\mathcal{A}}_{t}$ that contains $x$ for all $t\geq 0$ and $x\in X$. For all metric space $(X,\delta)$ and $\alpha>0$, define
$\gamma_{\alpha}\left(X,\delta\right)$ $\triangleq\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\Delta\left(A_{t}(x)\right)\,,$ (2)
where the infimum is taken over all admissible sequences of $X$.
Lemma 14.
For all $\alpha>0$, and metric spaces $(X,\delta_{1})$ and $(X,\delta_{2})$, if $\delta_{1}(x,y)\leq B\cdot\delta_{2}(x,y)$ for some $B>0$ and all $x,y\in X$, then $\gamma_{\alpha}\left(X,\delta_{1}\right)\leq B\cdot\gamma_{\alpha}\left(X,\delta_{2}\right)$.
Proof.
By (1) and (2),
$\gamma_{\alpha}\left(X,\delta_{1}\right)$ $=\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\left(\sup_{y,z\in A_{t}(x)}\delta_{1}\left(y,z\right)\right)$
$\leq B\cdot\inf\sup_{x\in X}\sum_{t\geq 0}2^{t/\alpha}\left(\sup_{y,z\in A_{t}(x)}\delta_{2}\left(y,z\right)\right)=B\cdot\gamma_{\alpha}\left(X,\delta_{2}\right)\,,$
where the inequality holds by $\delta_{1}(x,y)\leq B\cdot\delta_{2}(x,y)$ for all $x,y\in X$. ∎
Now, we are ready to prove the upper bound for $d=2$. Let $X=[0,1]^{2}$, $\delta(x,y)=\left\|{x-y}\right\|_{2}$ for all $x,y\in X$, and $\mathbb{D}$ be a $\beta$-smooth distribution on $X$. Our goal is to prove that
$\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\leq O\left(\sqrt{n\log n}\right)\,.$
Let ${\mathcal{C}}$ denote the class of $1$-lipschitz functions on $(X,\delta)$, i.e., functions $f$ such that $|f(x)-f(y)|\leq\delta(x,y)$ for all $x,y\in X$. For all $f\in{\mathcal{C}}$ and $S=\{s_{1},\ldots,s_{n}\}\subseteq X$, define
$Z_{f}(S)\triangleq\sum_{i=1}^{n}\left(f(s_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)\,.$
By [Tal92, Proposition $4.2.3$], for all $S=\{s_{1},\ldots,s_{n}\},R=\{r_{1},\ldots,r_{n}\}\subseteq X$,
$\mathrm{OPT}(S,R)$ $=\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(s_{i})-f(r_{i})\right)$
$\leq\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(s_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)+\sup_{f\in{\mathcal{C}}}\sum_{i=1}^{n}\left(f(r_{i})-\mathbb{E}_{x\sim\mathbb{D}}\left[f(x)\right]\right)$
$=\sup_{f\in{\mathcal{C}}}|Z_{f}(S)|+\sup_{f\in{\mathcal{C}}}|Z_{f}(R)|\,.$ (3)
Analogous to [Tal92, Corollary $4.3.5$], for every $v>0$, we can show that
$\mathbb{P}_{S\sim\mathbb{D}^{n}}[\left|Z_{f}(S)\right|\geq v]\leq 2\exp\left(-\min\left(\frac{v^{2}}{4n\left\|{f}\right\|_{2}^{\mathbb{D}}},\frac{v}{4\left\|{f}\right\|_{\infty}^{\mathbb{D}}}\right)\right)\,,$ (4)
where $\left\|{f}\right\|_{p}^{\mathbb{D}}$ denotes the norm of $f$ in $L_{p}$ with respect to the measure $\mu_{\mathbb{D}}$ induced by the distribution $\mathbb{D}$. The proof of (4) is identical to the proof of [Tal92, Corollary $4.3.5$] by applying Bernstein’s inequality and hence omitted here. Then, analogous to [Tal92, Theorem $4.3.6$], for every class ${\mathcal{F}}$ of functions on $[0,1]^{2}$ with $0\in{\mathcal{F}}$, we can show that
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{F}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{F}},d_{2}^{\mathbb{D}}\right)+\gamma_{1}\left({\mathcal{F}},d_{\infty}^{\mathbb{D}}\right)\right)\,,$ (5)
where $d_{2}^{\mathbb{D}}$ and $d_{\infty}^{\mathbb{D}}$ are the distances induced on ${\mathcal{F}}$ by the norms of $L_{2}$ and $L_{\infty}$ respectively, with respect to the measure $\mu_{\mathbb{D}}$ induced by $\mathbb{D}$ on $[0,1]^{2}$, and $\gamma_{\alpha}$ is defined in (2) for $\alpha>0$. The proof of (5) is identical to the proof of [Tal92, Theorem $4.3.6$] by applying (4) and [Tal92, Theorem $2.2.23$], and hence omitted here.
Let $d_{2}^{\mathbb{U}}$ and $d_{\infty}^{\mathbb{U}}$ be the distances induced on ${\mathcal{F}}$ by the norms of $L_{2}$ and $L_{\infty}$ respectively, with respect to the measure induced by the uniform distribution on $[0,1]^{2}$. For all $f,g\in\mathcal{F}$,
$d_{2}^{\mathbb{D}}(f,g)$ $=\left(\int_{X}(f(x)-g(x))^{2}\mathrm{d}\mu_{\mathbb{D}}(x)\right)^{1/2}\leq\sqrt{\beta}\left(\int_{X}(f(x)-g(x))^{2}\mathrm{d}x\right)^{1/2}$
$=\sqrt{\beta}\left(\mathbb{E}_{x\sim\mathbb{U}}\left[(f(x)-g(x))^{2}\right]\right)^{1/2}=\sqrt{\beta}\cdot d_{2}^{\mathbb{U}}(f,g)\,,$
where the inequality holds by the assumption that $\mathbb{D}$ is $\beta$-smooth. Also, note that
$d_{\infty}^{\mathbb{D}}(f,g)\leq d_{\infty}^{\mathbb{U}}(f,g)=\sup_{x\in X}|f(x)-g(x)|\,.$
Then, by Lemma 14 and (5), together with the above two displayed inequalities, given that $\beta>0$ is a constant, we have
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{F}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{F}},d_{2}^{\mathbb{U}}\right)+\gamma_{1}\left({\mathcal{F}},d_{\infty}^{\mathbb{U}}\right)\right)\,.$
Recall that ${\mathcal{C}}$ denotes the class of $1$-lipschitz functions on $(X,\delta)$. Then, we obtain
$\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{C}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{N}\gamma_{2}\left({\mathcal{C}},d_{2}^{\mathbb{U}}\right)+\gamma_{1}\left({\mathcal{C}},d_{\infty}^{\mathbb{U}}\right)\right)\leq O\left(\sqrt{n\log n}\right)\,,$ (6)
where the second inequality is implicitly shown in the proof of [Tal92, Theorem $4.3.2$]. Combining (3) and (6), it holds that
$\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\leq 2\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\sup_{f\in{\mathcal{C}}}\left|Z_{f}(S)\right|\right]\leq O\left(\sqrt{n\log n}\right)\,,$
which concludes the proof.
A.2 Proof of Lemma 3
For $d=1$, let $F(x)=\int_{0}^{x}\mu_{\mathbb{D}}(x)\mathrm{d}x$ for $x\in[0,1]$ be the cumulative density function of $\mathbb{D}$. Let $S=\{s_{1},\ldots,s_{n}\}\sim\mathbb{D}^{n}$ and $R=\{r_{1},\ldots,r_{n}\}\sim\mathbb{D}^{n}$ be two sets of random points drawn from $\mathbb{D}$. Define $P_{S}=\{F(s_{1}),\ldots,F(s_{n})\}$ and $P_{R}=\{F(r_{1}),\ldots,F(r_{n})\}$, and each point in $P_{S}\cup P_{R}$ follows uniform distribution on $[0,1]$. Since the expected cost of the optimal matching between two sets of $n$ uniformly distributed points on $[0,1]$ is $\Theta(\sqrt{n})$, we have
$\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]=\Theta\left(\sqrt{n}\right)\,.$ (7)
Note that for all $s,r\in[0,1]$,
$|F(s)-F(r)|=\left|\int_{r}^{s}f(x)\mathrm{d}x\right|\leq\left|\int_{r}^{s}\beta\mathrm{d}x\right|=\beta\cdot|s-r|\,.$ (8)
Let $M$ be the perfect matching between $S$ and $R$ with the minimum cost, and let $M^{\prime}$ be the perfect matching between $P_{S}$ and $P_{R}$ induced by $M$ such that $F(s)\in P_{S}$ is matched with $F(M(s))\in P_{R}$ for each $s\in S$. Hence, the cost of $M^{\prime}$ equals
$\sum_{s\in S}|F(s)-F(M(s))|\geq\mathrm{OPT}(P_{S},P_{R})\,.$ (9)
As a result,
$\mathrm{OPT}(S,R)=\sum_{s\in S}|s-M(s)|\geq\sum_{s\in S}\frac{1}{\beta}\cdot|F(s)-F(M(s))|\geq\frac{1}{\beta}\cdot\mathrm{OPT}(P_{S},P_{R})\,,$
where the equality holds by the definition of $M$, the first inequality holds by (8), and the second inequality holds by (9). Taking expectations on both sides, we obtain
$\mathrm{OPT}(n)=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq\frac{1}{\beta}\cdot\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(P_{S},P_{R})\right]\geq\Omega\left(\sqrt{n}\right)\,,$
where the last inequality holds by (7). This concludes the lower bound for $d=1$.
The lower bounds for $d\geq 2$ are proved by employing nearest-neighbor-distance, which is provided in the following lemma.
Lemma 15.
Assume $\mathbb{D}$ to be $\beta$-smooth for some constant $\beta>0$. For all $d\geq 1$ and $S\subseteq[0,1]^{d}$ with $|S|=n$,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(n^{-1/d}\right)\,.$
Proof.
Fix $S\subseteq[0,1]^{d}$ with $|S|=n$. For all $s\in S$ and $\ell>0$, let $B(s,\ell)=\{x\in[0,1]^{d}\mid\left\|{x-s}\right\|_{2}\leq\ell\}$ denote the ball restricted in $[0,1]^{d}$ with center $s$ and radius $\ell$, and we have
$\mathbb{P}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\leq\ell]=\int_{B(s,\ell)}\mu_{\mathbb{D}}(x)\mathrm{d}x\leq\int_{B(s,\ell)}\beta\mathrm{d}x\leq C\cdot\ell^{d}$
for some constant $C>0$. For all $s\in S$ and $r\in[0,1]^{d}$, let $O_{s,r}$ denote the event that $s\in\mathop{\arg\min}_{s^{\prime}\in S}\left\|{s^{\prime}-r}\right\|_{2}$ and $p_{s}=\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]$. For all $s\in S$ and $0<\ell\leq(p_{s}/C)^{1/d}$, it holds that
$\mathbb{P}_{r\sim\mathbb{D}}\left[O_{s,r}\land(\left\|{s-r}\right\|_{2}\geq\ell)\right]$ $=\mathbb{P}_{r\sim\mathbb{D}}\left[O_{s,r}\right]-\mathbb{P}_{r\sim\mathbb{D}}\left[O_{s,r}\land(\left\|{s-r}\right\|_{2}\leq\ell)\right]$
$\geq p_{s}-\mathbb{P}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\leq\ell\right]$
$\geq p_{s}-C\cdot\ell^{d}\,.$
As a result, for every $s\in S$,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\mid O_{s,r}\right]$ $\geq\int_{0}^{(p_{s}/C)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}\left[\left\|{s-r}\right\|_{2}\geq\ell\mid O_{s,r}\right]\mathrm{d}\ell$
$=\frac{1}{p_{s}}\int_{0}^{(p_{s}/C)^{1/d}}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}\land(\left\|{s-r}\right\|_{2}\geq\ell)]\mathrm{d}\ell$
$\geq\frac{1}{p_{s}}\int_{0}^{(p_{s}/C)^{1/d}}(p_{s}-C\cdot\ell^{d})\mathrm{d}\ell$
$=\frac{d}{d+1}\cdot\left(\frac{p_{s}}{C}\right)^{1/d}\,.$
Therefore,
$\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]=\sum_{s\in S}\mathbb{P}_{r\sim\mathbb{D}}[O_{s,r}]\cdot\mathbb{E}_{r\sim\mathbb{D}}[\left\|{s-r}\right\|_{2}\mid O_{s,r}]\geq\sum_{s\in S}\Omega\left(p_{s}^{1+\frac{1}{d}}\right)\geq\Omega\left(n^{-1/d}\right)\,,$
where the last inequality holds since $\sum_{s\in S}p_{s}=1$ and $p_{s}\geq 0$ for every $s\in S$. ∎
Now, we are ready to prove the lower bounds for $d\geq 2$. By Lemma 15,
$\mathrm{OPT}(n)$ $=\mathbb{E}_{S,R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]=\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\right]$
$\geq\mathbb{E}_{S\sim\mathbb{D}^{n}}\left[n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\right]\geq\Omega\left(n^{1-\frac{1}{d}}\right)\,,$
which concludes the proof.
A.3 Proof of Lemma 4
By Lemma 15 and the assumption that $|S|\leq O(n)$,
$\mathrm{OPT}(S,n)=\mathbb{E}_{R\sim\mathbb{D}^{n}}\left[\mathrm{OPT}(S,R)\right]\geq n\cdot\mathbb{E}_{r\sim\mathbb{D}}\left[\min_{s\in S}\left\|{s-r}\right\|_{2}\right]\geq\Omega\left(n^{1-\frac{1}{d}}\right)$
as desired.
Appendix B Postponed Proofs in Section 7.1
B.1 Proof of Lemma 12
We use ${\mathcal{A}}$ to denote the greedy subroutine in the MNP algorithm. The tie-breaking rule of ${\mathcal{A}}$ is local by definition, and it remains to show that it is also non-increasing. Recall that $N_{{\mathcal{A}}}(r\mid S)$ is defined as the set of available servers closest to the request $r$ upon its arrival when the set of initial servers is $S$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, for every $s\in S^{\prime}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}$ $=\frac{1}{|S^{\prime}|}\,,$
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}$ $=\frac{1}{|S^{\prime\prime}|}\,.$
Therefore,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,,$
which concludes that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing.
B.2 Proof of Lemma 13
We use ${\mathcal{A}}$ to denote the greedy subroutine in the Random Subtree algorithm. The tie-breaking rule of ${\mathcal{A}}$ is local by definition, and it remains to show that it is also non-increasing. For each node $x$ in the HST, denote $A_{\ell}(x)$ as the $\ell$-th ancestor of $x$ and $T_{\ell}(x)$ as the subtree rooted at $A_{\ell}(x)$. That is, $A_{0}(x)$ is $x$, $A_{1}(x)$ is the parent of $x$, and so on. Fix a request $r$. For all sets of servers $S^{\prime},S^{\prime\prime}$ with $S^{\prime}\subseteq S^{\prime\prime}$, let $\ell^{\prime}$ satisfy $A_{\ell^{\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$, and let $\ell^{\prime\prime}$ satisfy $A_{\ell^{\prime\prime}}(r)=A(r\mid S)$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. Note that $\ell^{\prime}$ and $\ell^{\prime\prime}$ only depend on $S^{\prime}$ and $S^{\prime\prime}$, respectively.
We first show that $\ell^{\prime}=\ell^{\prime\prime}$ by deriving a contradiction when assuming $\ell^{\prime}\neq\ell^{\prime\prime}$. We only analyze the case where $\ell^{\prime}<\ell^{\prime\prime}$, and the case where $\ell^{\prime}>\ell^{\prime\prime}$ can be proved analogously. When $\ell^{\prime}<\ell^{\prime\prime}$, all servers in $S^{\prime}$ are located in $T_{\ell^{\prime}}(r)$, and all servers in $S^{\prime\prime}$ are located in $T_{\ell^{\prime\prime}}(r)$. Moreover, by the definition of $\ell^{\prime\prime}$, no server in $S^{\prime\prime}$ is located in $T_{\ell^{\prime\prime}-1}(r)$. Since $T_{\ell^{\prime}}(r)$ is a subtree in $T_{\ell^{\prime\prime}-1}(r)$, we have $S^{\prime}\cap S^{\prime\prime}=\emptyset$, contradicting the assumption that $S^{\prime}\subseteq S^{\prime\prime}$.
Fix a server $s\in S^{\prime}$. For every $i\in[\ell^{\prime}]$, denote $w_{i}^{\prime}$ as the number of subtrees of $A_{i}(s)$ containing at least one available server upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Moreover, define $w_{0}^{\prime}$ as the number of available servers collocated with $s$ (including $s$) upon the arrival of $r$ when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime}$. Recall that $\ell^{\prime}=\ell^{\prime\prime}$, and define $w_{i}^{\prime\prime}$ for all $i\in\{0,1,\ldots,\ell^{\prime}\}$ similarly when $N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}$. By the definition of the tie-breaking rule of ${\mathcal{A}}$,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}$ $=\prod_{i=0}^{\ell^{\prime}}\frac{1}{w_{i}^{\prime}}\,,$
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}$ $=\prod_{i=0}^{\ell^{\prime}}\frac{1}{w_{i}^{\prime\prime}}\,.$
Since $S^{\prime}\subseteq S^{\prime\prime}$, we have $w_{i}^{\prime}\leq w_{i}^{\prime\prime}$ for all $i\in\{0,1,\ldots,\ell^{\prime}\}$. Therefore,
$\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime}\right\}\geq\mathbb{P}\left\{{\mathcal{A}}\text{ matches }r\text{ to }s\mid N_{{\mathcal{A}}}(r\mid S)=S^{\prime\prime}\right\}\,,$
concluding that the tie-breaking rule of ${\mathcal{A}}$ is non-increasing.
Appendix C Stochastic Setting: Cost Functions Beyond Triangle Inequality
In previous sections, we consider metric space where the cost function satisfies the exact triangle inequality. In this section, we showcase the generality of our reduction for the stochastic setting by proving positive results when the cost function only approximately satisfies the triangle inequality. Specifically, we say that a cost function $\delta$ on $X$ satisfies the $\eta$-approximate triangle inequality for $\eta>0$ if for all $x,y,z\in X$,
$\delta(x,z)\leq\eta\cdot(\delta(x,y)+\delta(y,z))\,.$
We say that a metric space satisfies the $\eta$-approximate triangle inequality if the corresponding cost function satisfies the $\eta$-approximate triangle inequality.
We first generalize Theorem 2 to the case where the metric space only satisfies the approximate triangle inequality.
Theorem 10.
Given a distribution $\mathbb{D}$ on a metric space $(X,\delta)$ satisfying the $\eta$-approximate triangle inequality for $\eta>0$, assume the existence of an algorithm ${\mathcal{A}}$ such that $\mathrm{cost}_{{\mathcal{A}}}(n)\leq\alpha\cdot\mathrm{OPT}(n)$. Then, an algorithm ${\mathcal{A}}^{\prime}$ exists such that $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq\eta\cdot(\mathrm{OPT}(S,n)+\mathrm{cost}_{{\mathcal{A}}}(n))$ for every set of servers $S$ with $|S|\geq n$. If we further have $|S|=n$, then $\mathrm{cost}_{{\mathcal{A}}^{\prime}}(S,n)\leq(2\alpha\eta^{2}+\eta)\cdot\mathrm{OPT}(S,n)$.
The proof of Theorem 10 can be obtained by replacing every usage of the triangle inequality by the approximate triangle inequality in the proofs of Theorem 1 and Theorem 2 and hence is omitted.
Next, we consider the case where $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $p\geq 1$, [^6] Here, $p$ can be a real number. which we will prove that satisfies the $2^{p-1}$-approximate triangle inequality, and assume $\mathbb{D}$ to be uniform distribution. We show that in balanced markets with adversarial servers and stochastic requests, an $O(1)$-competitive algorithm exists for $d\geq 1$ and $p\in[1,\max\{d,2\})$.
Theorem 11.
Assume that $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $d\geq 1$ and $p\in[1,\max\{d,2\})$. Supposing that $m=n$ and $\mathbb{D}$ is uniform distribution, an $O(1)$-competitive algorithm exists with adversarial servers and stochastic requests.
We remark here that competitive ratio guarantees for other combinations of $d$ and $p$, more general distributions, and unbalanced markets can be similarly achieved as in Section 4 and Section 5. Nevertheless, we didn’t pursue this direction since parts of the resulting competitive ratios seem suboptimal, i.e., super-constant, and no new algorithmic insight emerges.
C.1 Proof of Theorem 11
We start by showing that for all $p\geq 1$, the cost function $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ satisfies the $2^{p-1}$-approximate triangle inequality by proving a more general statement: Assuming $\delta^{\prime}$ to be a cost function satisfying the triangle inequality, the cost function defined as $\delta(x,y)\triangleq\delta^{\prime}(x,y)^{p}$ satisfies the $2^{p-1}$-approximate triangle inequality.
Lemma 16.
For every $p\geq 1$, given a cost function $\delta^{\prime}$ on $X$ satisfying the triangle inequality, the cost function on $X$ defined as $\delta(x,y)\triangleq\delta^{\prime}(x,y)^{p}$ for all $x,y\in X$ satisfies the $2^{p-1}$-approximate triangle inequality.
Proof.
For all $x,y,z\in X$,
$\delta(x,z)$ $=\delta^{\prime}(x,z)^{p}\leq(\delta^{\prime}(x,y)+\delta^{\prime}(y,z))^{p}=\left(\frac{\delta^{\prime}(x,y)}{2}+\frac{\delta^{\prime}(y,z)}{2}\right)^{p}\cdot 2^{p}$
$\leq\left(\frac{\delta^{\prime}(x,y)^{p}}{2}+\frac{\delta^{\prime}(y,z)^{p}}{2}\right)\cdot 2^{p}=(\delta(x,y)+\delta(y,z))\cdot 2^{p-1}\,,$
where the first inequality holds by the triangle inequality and the second inequality holds by the convexity of $g(x)=x^{p}$ with $p\geq 1$. ∎
Next, we present the following estimation of $\mathrm{OPT}(n)$.
Lemma 17 ([Led19]).
Let $X=[0,1]^{d}$ and $\delta(x,y)=\left\|{x-y}\right\|_{2}^{p}$ for $d,p\geq 1$. Assuming ${\mathbb{D}}$ to be uniform distribution, it holds that
$\mathrm{OPT}(n)=\begin{cases}\Theta(n^{1-\frac{p}{2}})&d=1\\ \Theta(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}})&d=2\\ \Theta(n^{1-\frac{p}{d}})&d\geq 3\end{cases}\,.$
Now, we are ready to prove Theorem 11. Let ${\mathcal{A}}$ be the algorithm as per Lemma 5, and we analyze $\mathrm{cost}_{{\mathcal{A}}}(n)$ for $d=1$, $d=2$, and $d\geq 3$ separately by applying Lemma 17. For $d=1$ and $p\in[1,2)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{2}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
For $d=2$ and $p\in[1,2)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{2}}(\log t)^{\frac{p}{2}}\right)=O\left((\log n)^{\frac{p}{2}}\int_{1}^{n}x^{-\frac{p}{2}}\mathrm{d}x\right)$
$=O\left(n^{1-\frac{p}{2}}(\log n)^{\frac{p}{2}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
For $d\geq 3$ and $p\in[1,d)$,
$\mathrm{cost}_{{\mathcal{A}}}(n)=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}t^{-\frac{p}{d}}\right)=O\left(\int_{1}^{n}x^{-\frac{p}{d}}\mathrm{d}x\right)=O\left(n^{1-\frac{p}{d}}\right)=O\left(\mathrm{OPT}(n)\right)\,.$
As a result, ${\mathcal{A}}$ admits an $O(1)$ competitive ratio when both servers and requests are uniformly distributed. By Lemma 16, $\delta$ satisfies the $2^{p-1}$-approximate triangle inequality. Combined with Theorem 10, we conclude that an $O(1)$-competitive algorithm exists in balanced markets with adversarial servers and uniformly distributed requests.
Appendix D Stochastic Setting: An Alternative Algorithm for Tree Metrics
In this section, we consider balanced markets and apply our approach to tree metrics with general distributions. In particular, a tree metric is defined by shortest path distances in a tree $T=(V,E)$. We will prove that the $\mathsf{SOAR}$ algorithm admits an $O(1)$ competitive ratio with all servers and requests independently drawn from $\mathbb{D}$, which implies the existence of an $O(1)$-competitive algorithm with servers adversarially chosen and requests independently drawn from $\mathbb{D}$ by Theorem 2. This gives an alternative $O(1)$-competitive algorithm in balanced markets with tree metrics and general distributions, recovering the result by [GGPW19].
We first present our main theorem in this section.
Theorem 12.
Suppose that $m=n$ and consider tree metrics. For an arbitrary distribution $\mathbb{D}$, an $O(1)$-competitive algorithm exists when servers are adversarially chosen and requests are independently drawn from $\mathbb{D}$.
To prove Theorem 12, it is known that in balanced markets with adversarial servers and stochastic requests following an arbitrary distribution, only a constant factor in the competitive ratio is lost by assuming that each request is uniformly distributed over server locations. That is, for every server $s\in S$, the request is collocated with $s$ with probability $1/n$.
Lemma 18 ([GGPW19]).
Suppose that $m=n$ and servers are adversarially chosen. For an arbitrary metric, if an $\alpha$-competitive algorithm exists when requests are uniformly distributed over server locations, then we can construct a $(2\alpha+1)$-competitive algorithm when requests follow an arbitrary distribution.
Given Lemma 18, it suffices to consider distribution $\mathbb{D}$ of the following form: Given a tree metric $T=(V,E)$, each node $v\in V$ is associated with an integer $a_{v}\geq 0$ such that $\sum_{v\in V}a_{v}=n$ and $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$. For each edge $e\in E$, let $d_{e}$ denote the length of $e$, and let $T_{1}(e)$ and $T_{2}(e)$ be two subtrees of $T$ obtained by deleting $e$. For each $e\in E$, define $n_{e}=\sum_{v\in T_{1}(e)}a_{v}$. Given such a distribution $\mathbb{D}$, we estimate $\mathrm{OPT}(t)$ in the following lemma.
Lemma 19.
For each $t\in[n]$, $\mathrm{OPT}(t)=\Theta\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)$.
Proof.
Fix $t\in[n]$. For each $e\in E$, let $P_{e},Q_{e}\sim{\rm Binom}(t,n_{e}/n)$ be random variables denoting the number of servers located in $T_{1}(e)$ and the number of requests located in $T_{1}(e)$, respectively. Given a realization of $(P_{e},Q_{e})_{e\in E}$, in the perfect matching between servers and requests with the minimum cost, for each $e\in E$, exactly $|P_{e}-Q_{e}|$ matched server-request pairs will utilize $e$, incurring a cost of $d_{e}\cdot|P_{e}-Q_{e}|$. Hence,
$\mathrm{OPT}(t)=\mathbb{E}\left[\sum_{e\in E}d_{e}\cdot|P_{e}-Q_{e}|\right]=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\,.$
We first lower bound $\mathrm{OPT}(t)$, which relies on the following probabilistic bound.
Claim 1 ([BK13]).
Let $Z\sim{\rm Binom}(n,p)$, with $n\geq 2$ and $p\in[1/n,1-1/n]$. Then, we have
$\mathbb{E}[|Z-\mathbb{E}Z|]\geq\mathrm{std}(Z)/\sqrt{2}\,,$
where $\mathrm{std}(Z)$ denotes the standard deviation of $Z$.
We lower bound $\mathrm{OPT}(t)$ by
$\mathrm{OPT}(t)$ $=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\geq\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-\mathbb{E}P_{e}|]$
$\geq\frac{1}{\sqrt{2}}\sum_{e\in E}d_{e}\cdot\mathrm{std}(P_{e})=\Omega\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)\,,$
where the first inequality holds since the function $f(z)=\mathbb{E}[|P_{e}-z|]$ reaches its minimum when $z=\mathbb{E}P_{e}$, and the second inequality holds by Claim 1. Moreover, $\mathrm{OPT}(t)$ is upper bounded by
$\mathrm{OPT}(t)$ $=\sum_{e\in E}d_{e}\cdot\mathbb{E}[|P_{e}-Q_{e}|]\leq\sum_{e\in E}d_{e}\left(\mathbb{E}\left[\left|P_{e}-\mathbb{E}P_{e}\right|\right]+\mathbb{E}\left[\left|Q_{e}-\mathbb{E}Q_{e}\right|\right]\right)$
$\leq\sum_{e\in E}d_{e}(\mathrm{std}(P_{e})+\mathrm{std}(Q_{e}))=O\left(\frac{\sqrt{t}}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)\,,$
where the first inequality holds by $\mathbb{E}P_{e}=\mathbb{E}Q_{e}$ for every $e\in E$. ∎
Now, we are ready to prove Theorem 12.
Proof of Theorem 12.
According to Lemma 18, we can assume requests to be uniformly distributed over server locations by losing a constant factor in the competitive ratio. In other words, let $a_{v}$ be the number of servers located at $v\in V$, the distribution $\mathbb{D}$ satisfies $\mathbb{P}_{r\sim\mathbb{D}}[r=v]=a_{v}/n$ for every $v\in V$. By Theorem 2, it suffices to give an $O(1)$-competitive algorithm when all servers and requests are independently drawn from $\mathbb{D}$.
Let ${\mathcal{A}}$ be the algorithm as per Lemma 5. Recall that $\mathrm{OPT}(t)$ is estimated by Lemma 19. When all servers and requests are independently drawn from $\mathbb{D}$, the expected cost of ${\mathcal{A}}$ equals
$\mathrm{cost}_{{\mathcal{A}}}(n)$ $=\sum_{t=1}^{n}\frac{1}{t}\cdot\mathrm{OPT}(t)=O\left(\sum_{t=1}^{n}\frac{1}{n\sqrt{t}}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)$
$=O\left(\frac{1}{n}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\sum_{t=1}^{n}\frac{1}{\sqrt{t}}\right)=O\left(\frac{1}{\sqrt{n}}\sum_{e\in E}d_{e}\sqrt{n_{e}(n-n_{e})}\right)=O\left(\mathrm{OPT}(n)\right)\,,$
concluding that ${\mathcal{A}}$ is $O(1)$-competitive. ∎